### v1

Abstract
Nonsmooth nonconvex-concave minimax problems have attracted significant attention due to their wide applications in many fields. In this paper, we consider a class of nonsmooth nonconvex-concave minimax problems on Riemannian manifolds. Owing to the nonsmoothness of the objective function, existing minimax manifold optimization methods cannot be directly applied to solve this problem. We propose a manifold proximal gradient descent ascent (MPGDA) algorithm for solving the problem. At each iteration, the proposed algorithm alternately performs one or multiple manifold proximal gradient descent steps and a proximal ascent step. We prove that the MPGDA algorithm can find an $\varepsilon$-game-stationary point and an $\varepsilon$-optimization-stationary point of the considered problem within $\mathcal{O}(\varepsilon^{-3})$ iterations. Numerical experiments on fair sparse PCA and sparse spectral clustering are conducted to demonstrate the advantages of the MPGDA algorithm.
Keywords:
nonconvex-concave minimax problem nonsmooth manifold optimization manifold proximal gradient descent ascent iteration complexity
MSC:
90C26 90C30 65K05
1 Introduction
In this paper, we consider a class of nonsmooth nonconvex-concave minimax optimization problems formulated as follows:
$\min_{x\in\mathcal{M}}\max_{y\in S}F(x,y):=f(x,y)+h(x)-g(y),$ (1)
where $\mathcal{M}$ is a Riemannian manifold embedded in a finite-dimensional Euclidean space $\mathbb{E}_{1}$, and $S$ is a convex set in another finite-dimensioanl Euclidean space $\mathbb{E}_{2}$, $f:\mathbb{E}_{1}\times\mathbb{E}_{2}\to\mathbb{R}$ is a continuously differentiable function and is concave with respect to y, $h:\mathbb{E}_{1}\to\mathbb{R}$ and $g:\mathbb{E}_{2}\to\mathbb{R}$ are some proper closed convex (possibly nonsmooth) functions. Throughout this paper, we adopt the following assumptions for problem (1) .
{assumption}
$(i)$ The manifold $\mathcal{M}$ is compact.
$(ii)$ The function $f$ has Lipschitz continuous gradient with respect to $x$, i.e., there exists a Lipschitz constant $L_{x}>0$ such that
$\left\|\nabla_{x}f(x_{1},y)-\nabla_{x}f(x_{2},y)\right\|\leq L_{x}\left\|x_{1}-x_{2}\right\|$
for any $x_{1},x_{2}\in\mathcal{M}$ and $y\in S$.
$(iii)$ The function $\nabla_{y}f(x,y)$ is Lipschitz continuous with respect to $(x,y)$, i.e., there exists a Lipschitz constant $L_{y}>0$ such that
$\left\|\nabla_{y}f(x_{1},y_{1})-\nabla_{y}f(x_{2},y_{2})\right\|^{2}\leq L_{y}^{2}(\left\|x_{1}-x_{2}\right\|^{2}+\left\|y_{1}-y_{2}\right\|^{2})$
for any $(x_{1},y_{1})$, $(x_{2},y_{2})\in\mathcal{M}\times S$.
$(iv)$ The function $h$ is $L_{h}$ Lipschitz continuous.
Problem (1) has many important applications in machine learning and signal processing. Below, we present two representative examples and refer the interested readers to chenProximalGradientMethod2020 ; BriefIntroductiontoManifoldOptimization for more examples.
Example 1.
Fair Sparse Principle Component Analysis (FSPCA). The classical PCA Hotelling1933Analysis is one of the most widely used dimensional reduction techniques, which minimizes the total reconstruction error over the whole samples. However, in real-life applications, the data samples come from diverse classes and PCA may exhibit a higher reconstruction error for certain classes than the others. To address this disparity, the fair PCA is proposed samadiPriceFairPCA . Suppose the $m$ data samples belong to $n$ classes, and each class $i$ corresponds to $A_{i}\in\mathbb{R}^{m_{i}\times d}$ with $d$ being the dimension of the samples, $m_{i}$ being the number of samples in the $i$-th class, and $\sum_{i=1}^{n}m_{i}=m$. Then the fair PCA can be formulated as:
$\min_{X\in\mathrm{St}(d,r)}\max_{i=1,\dots,n}-\mathrm{Tr}(X^{\top}A_{i}^{\top}A_{i}X),$ (2)
where $r<d$ is the number of principle components, $\mathrm{Tr}(\cdot)$ represents the trace and $\mathrm{St}(d,r):=\{X\in\mathbb{R}^{d\times r}|X^{\top}X=I_{r}\}$ with $I_{r}$ denoting the $r\times r$ identity matrix. Recently, the fair sparse PCA is further proposed to promote sparsity in the principle components babuFairPrincipalComponent2023 as follows:
$\min_{X\in\mathrm{St}(d,r)}\max_{i=1,\dots,n}-\mathrm{Tr}(X^{\top}A_{i}^{\top}A_{i}X)+\mu\left\|X\right\|_{1},$ (3)
where $\mu>0$ is a weighting parameter and $\left\|X\right\|_{1}:=\sum_{i,j}\left|x_{ij}\right|$ is the $\ell_{1}$-norm of the matrix $X$. In view of (Beck2017First, , Example 4.10), problem (3) can be equivalently reformulated into
$\min_{X\in\mathrm{St}(d,r)}\max_{y\in\Delta_{n}}-\sum_{i=1}^{n}y_{i}\mathrm{Tr}(X^{\top}A_{i}^{\top}A_{i}X)+\mu\left\|X\right\|_{1}.$ (4)
where $\Delta_{n}:=\{y|\sum_{i=1}^{n}y_{i}=1,y_{i}\geq 0,i=1,\dots,n\}$ is the standard simplex in $\mathbb{R}^{n}$. Note that problem (4) is an instance of problem (1) with $\mathcal{M}$ being the Stiefel manifold $\mathrm{St}(d,r)$, $S$ being the standard simplex $\Delta_{n}$, $h(X)=\mu\|X\|_{1}$ and $f(X,y)=-\sum_{i=1}^{n}y_{i}\mathrm{Tr}(X^{\top}A_{i}^{\top}A_{i}X)$.
Example 2.
Sparse Spectral Clustering (SSC). The SSC problem aims to partition $N$ data samples into $p$ groups such that similar data points are clustered together. In spectral clustering, a symmetric affinity matrix $W=[W_{ij}]_{N\times N}$ is constructed, where $W_{ij}\geq 0$ quantifies the pairwise similarity between samples $a_{i}$ and $a_{j}$. To promote sparsity and interpretability, the sparse spectral clustering framework is introduced in 2016ConvexSparseSpectralClustering , as follows
$\min_{X\in\mathrm{St}(N,p)}\langle L,XX^{\top}\rangle+\mu\|XX^{\top}\|_{1},$ (5)
where $\mu>0$ is the regularization parameter and $L=I_{N}-S^{-1/2}WS^{-1/2}$ is the normalized Laplacian matrix with $S^{1/2}$ being the diagonal matrix with diagonal elements $\sqrt{s_{1}},\sqrt{s_{2}},\dots,\sqrt{s_{n}}$ and $s_{i}=\sum_{j}W_{ij}$. It is not hard to see that problem (5) can be equivalently rewritten as:
$\min_{X\in\mathrm{St}(N,p),~{}Z\in\mathbb{R}^{N\times N}}\langle L,XX^{\top}\rangle+\mu\|Z\|_{1}+\mu\|XX^{\top}-Z\|_{1}.$ (6)
By the dual representation of the $\ell_{1}$-norm, i.e., $\|\cdot\|_{1}=\max_{\|Y\|_{\infty}\leq 1}\langle Y,\cdot\rangle$, problem (6) can be further equivalently reformulated as:
$\min_{X\in\mathrm{St}(N,p),~{}Z\in\mathbb{R}^{N\times N}}\max_{\|Y\|_{\infty}\leq\mu}\langle L,XX^{\top}\rangle+\mu\|Z\|_{1}+\langle Y,XX^{\top}-Z\rangle.$ (7)
Again, problem (7) is an instance of problem (1) with $\mathcal{M}$ being the product manifold $\mathrm{St}(N,p)\times\mathbb{R}^{N\times N}$, $S=\{Y\in\mathbb{R}^{N\times N}\mid\|Y\|_{\infty}\leq\mu\}$, $h(X,Z)=\mu\|Z\|_{1}$ and $f(X,Z,Y)=\langle L,XX^{\top}\rangle+\langle Y,XX^{\top}-Z\rangle$.
In recent years, nonconvex-concave (NC-C) minimax problems have received a tremendous amount of attention as they have found broad applications in diverse areas. Many studies focus on NC-C problems in the setting where both the constraints on $x$ and $y$ are closed convex sets and numerous methods have been developed, see, for example, linGradientDescentAscent ; linOptimalAlgorithmsMinimax ; nouiehedSolvingClassNonConvex ; ostrovskiiEfficientSearchFirstOrder2021 ; panEfficientAlgorithmNonconvexlinear2021a ; 2022Zeroth ; thekumparampilEfficientAlgorithmsSmooth ; xuUnifiedSingleloopAlternating2023 ; yangFasterSingleloopAlgorithms ; zhangSingleLoopSmoothedGradient . In contrast, there are relatively few works studying NC-C minimax problems in the setting where the constraint on $x$ is a Riemannian submanifold and the constraint on $y$ is a closed convex set. Among these works, Huang and Gao huangGradientDescentAscent2023 propose a Riemannian gradient descent-ascent algorithm for a smooth nonconvex-strongly concave (NC-SC) minimax problems and establish its iteration complexity of $\mathcal{O}(\varepsilon^{-2})$ for achieving $\varepsilon$-stationarity. Later, Xu et al. xuEfficientAlternatingRiemannian2024 present an alternating Riemannian/projected gradient descent-ascent (ARPGDA) algorithm with an iteration complexity of $\mathcal{O}(\varepsilon^{-3})$ for finding an $\varepsilon$-stationary point of smooth Riemannian nonconvex-linear (NC-L) minimax problem. Very recently, for problem (1) with additional assumptions that $h$ is smooth (but not necessarily convex) and $f(x,y)$ is linear with respect to $y$ for fixed $x$, the work xuRiemannianAlternatingDescent2024 extend the technique in and develop algorithms that can reach $\varepsilon$-stationary points within $\mathcal{O}(\varepsilon^{-3})$ iterations. It is worth noting that none of the methods in the aforementioned literature huangGradientDescentAscent2023 ; xuRiemannianAlternatingDescent2024 ; xuEfficientAlternatingRiemannian2024 can directly tackle problem (1) due to the following two reasons: (i) the convex function $h$ may be nonsmooth, and (ii) the function $f(x,y)$ is neither linear nor strongly concave with respect to $y$ for fixed $x$. Based on the needs of applications and the above background, it is highly desirable to develop efficient methods with theoretical convergence guarantees for solving the Riemannian nonsmooth NC-C minimax problem (1).
The main contributions of this paper are summarized below. First, we introduce the notions of optimization stationarity point as well as game stationarity for problem (1) and carefully study their relationships. Second, we propose a manifold proximal gradient descent ascent method for solving problem (1). At each iteration, this method performs one or multiple manifold proximal gradient descent steps to update $x$ and one proximal ascent step to update $y$. Third, we prove that the proposed method can find an $\varepsilon$-game-stationary point and an $\varepsilon$-optimization-stationary point within $O(\varepsilon^{-3})$ iterations. This iteration complexity result matches the state-of-the-art complexity analysis for single-loop algorithms in solving general nonsmooth NC-C optimization problems, where the constraint on $x$ is a compact convex set rather than a Riemannian manifold. Finally, we present numerical results on FSPCA and SSC to demonstrate the effectiveness of our proposed method.
The rest of the paper is organized as follows. Section 2 presents preliminaries that will be used in this paper. Section 3 discusses the optimality condition of problem (1). Section 4 introduces the proposed MPGDA method and establish its iteration complexity. Section 5 reports the numerical experiments.
2 Preliminaries
[…]
2.1 Riemannian submanifold
We first review the definition of an embedded submanifold in a Euclidean space $\mathbb{E}$. We say that function $\psi:U\subset\mathbb{E}\to\mathbb{R}^{k}$ is smooth if it is infinitely differentiable on $U$. Let $\mathrm{D}\psi(x)$ denote its differential at $x$.
Definition 1.
(boumal2023intromanifolds, , Definition 3.10) Let $\mathbb{E}$ be a Euclidean space of dimension $d$. A nonempty subset $\mathcal{M}$ of $\mathbb{E}$ is called an embedded submanifold of $\mathbb{E}$ of dimension $n$ if either $n=d$ and $\mathcal{M}$ is open in $\mathbb{E}$ or $n=d-k$ for some $k\geq 1$ and, for each $x\in\mathcal{M}$, there exists a neighborhood $U$ of $x$ in $\mathbb{E}$ and a smooth function $\psi:U\to\mathbb{R}^{k}$ such that:
  1. 1.
    If $y$ is in $U$, then $\psi(y)=0$ if and only if $y\in\mathcal{M}$; and
  2. 2.
    $\mathrm{rank}(\mathrm{D}\psi(x))=k$.
Such a function $\psi$ is called a local defining function for $\mathcal{M}$ at $x$.
Let $\mathcal{M}$ be an embedded submanifold of $\mathbb{E}$. The tangent space $T_{x}\mathcal{M}$ to $\mathcal{M}$ at $x\in\mathcal{M}$ is identified with $\mathbb{E}$ when $\mathcal{M}$ is open in $\mathbb{E}$; otherwise, $T_{x}\mathcal{M}=\mathrm{ker}(\mathrm{D}\psi(x))$ , where $\psi$ is a local defining function for $\mathcal{M}$ at $x$ (boumal2023intromanifolds, , Theorem 3.15). The embedded submanifold $\mathcal{M}$ is called a Riemannian submanifold of $\mathbb{E}$, when equipped with the Riemannian metric induced by the ambient space $\mathbb{E}$. Namely, for each $x\in\mathcal{M}$, the inner product on $T_{x}\mathcal{M}$ is defined as $\langle u,v\rangle_{x}:=\langle u,v\rangle$ for $u,~{}v\in T_{x}\mathcal{M}$. The normal space $N_{x}\mathcal{M}$ of a Riemannian submanifold $\mathcal{M}$ of $\mathbb{E}$ is the orthogonal complement of the tangent space $T_{x}\mathcal{M}$ in $\mathbb{E}$ (boumal2023intromanifolds, , Definition 5.48).
The tangent bundle of $\mathcal{M}$ is the disjoint union of its tangent spaces : $T\mathcal{M}:=\{(x,v)\mid x\in\mathcal{M},~{}v\in T_{x}\mathcal{M}\}$. A retraction on a manifold $\mathcal{M}$ is a smooth mapping $R$ from the tangent bundle $T\mathcal{M}$ onto $\mathcal{M}$ such that, for any $(x,v)\in T_{x}\mathcal{M}$, its restriction $R_{x}:T_{x}\mathcal{M}\to\mathcal{M}$ satisfies (i) $R_{x}(0)=x$ and (ii) $\mathrm{D}R_{x}(0)[v]=v$ (boumal2023intromanifolds, , Definition 3.47). The concept of a retraction offers a practical link between manifold and its tangent bundle. The following properties of retractions are needed for convergence analysis in this paper.
Proposition 1.
[…]
$\left\|R_{x}(v)-x\right\|$ $\leq M_{1}\left\|v\right\|,$
$\left\|R_{x}(v)-(x+v)\right\|$ $\leq M_{2}\left\|v\right\|^{2}.$
2.2 Generalized subdifferential
For a fucntion $\varphi:\mathbb{E}\to\mathbb{R}$, the usual one-sided directional derivative of $\varphi$ at $x\in\mathbb{E}$ along the direction $v\in\mathbb{E}$ is defined by
$\varphi^{\prime}(x;v):=\lim_{t\downarrow 0}\frac{\varphi(x+tv)-\varphi(x)}{t}.$
In contrast to the usual one-sided directional derivative, the (Clarke) generalized directional derivative dose not presuppose the existence of any limit (1983Optimization, , Section 2.1). Specifically, the (Clarke) generalized directional derivative of $\varphi$ at $x\in\mathbb{E}$ along the direction $v\in\mathbb{E}$ is defined as
$\varphi^{\circ}(x;v):=\lim\sup_{\begin{smallmatrix}y\to x\\ t\downarrow 0\end{smallmatrix}}\frac{\varphi(y+tv)-\varphi(y)}{t},$
and the (Clarke) generalized subdifferential of $\varphi$ at $x$ is defined as
$\partial\varphi(x):=\{z\in\mathbb{E}\mid\langle z,v\rangle\leq\varphi^{\circ}(x;v),~{}\forall v\in\mathbb{E}\}.$
The function $\varphi$ is said to be regular at $x$, provided that for all $v\in\mathbb{E}$, the usual one-sided directional derivative $\varphi^{\prime}(x;v)$ exists and $\varphi^{\prime}(x;v)=\varphi^{\circ}(x;v)$. If $\varphi_{1}$ and $\varphi_{2}$ are regular at $x$, then $\partial(\varphi_{1}+\varphi_{2})(x)=\partial\varphi_{1}(x)+\partial\varphi_{2}(x)$ (1983Optimization, , Definition 2.3.4). If $\varphi$ is a continuously differentiable function or a real-valued convex function, then it is regualr at any $x\in\mathbb{E}$ (1983Optimization, , Proposition 2.3.6). In addition, the indication function $\partial\delta_{\mathcal{M}}(x)$ of a Remannian submanifold $\mathcal{M}$ in $\mathbb{E}$ is regular at any $x\in\mathcal{M}$ (1983Optimization, , Theorem 2.4.7) and $\partial\delta_{\mathcal{M}}(x)=N_{x}\mathcal{M}$ (1983Optimization, , Proposition 2.4.12).
3 Optimality Conditions
In this section, we introduce two important concepts of stationarity for problem (1), namely, the optimization stationarity and game stationarity, and examine their relationships. To this end, we first definite the value function $\Phi:\mathbb{E}\to\mathbb{R}$ by
$\Phi(x):=\max_{y\in S}f(x,y)-g(y).$ (8)
Then the minimax problem (1) can be equivalently reformulated into
$\min_{x\in\mathcal{M}}\Phi(x)+h(x).$ (9)
The next proposition concerns the value function $\Phi$. Recalled that, a function $\varphi:\mathbb{E}\to\mathbb{R}$ is said to be $\ell$-weakly convex if $\varphi(\cdot)+(\ell/2)\|\cdot\|^{2}$ is convex.
Proposition 2.
Let $y_{x}\in\arg\max_{y\in S}\{f(x,y)-g(y)\}$. The value function $\Phi$ is $L_{x}$-weakly convex and for all $x\in\mathcal{M}$, it holds that $\nabla_{x}f(x,y_{x})\in\partial\Phi(x)$. Moreover, if $f(x,y)-g(y)$ is strongly concave in $y$, then $\Phi$ is differentiable and $\nabla\Phi(x)=\nabla_{x}f(x,y_{x})$.
Proof.
Because $f(x,y)$ is $L_{x}$-smooth respect to $x$, $f(x,y)+\frac{L_{x}}{2}\left\|x\right\|^{2}-g(y)$ is convex in $x$ for any given $y\in S$. Thus, $\Phi(x)+\frac{L_{x}}{2}\left\|x\right\|^{2}=\max_{y\in S}f(x,y)+\frac{L_{x}}{2}\left\|x\right\|^{2}-g(y)$ is convex. Therefore, $\Phi(x)$ is $L_{x}$-weakly convex. Since $S$ is compact, according to Danskin’s theorem (bernhardTheoremDanskinApplication1995, , Theorem D1), $\nabla_{x}f(x,y_{x})\in\partial\Phi(x)$, and in particular, $\nabla\Phi(x)=\nabla_{x}f(x,y_{x})$ when $f(x,y)-g(y)$ is strongly concave in $y$. ∎
In view of Proposition 2 and (1983Optimization, , Proposition 2.4.3), we see that any local minimizer $x^{*}$ of problem (9) must satisfy
$0\in\partial\Phi(x^{*})+\partial h(x^{*})+\partial\delta_{\mathcal{M}}(x^{*}).$ (10)
This naturally leads to the following definition of optimization stationarity.
Definition 2 (Optimization stationarity).
Consider problem (1) and the value function $\Phi$ defined in (8).
  1. (i)
    We say $x^{*}$ is a optimization-stationary point of problem (1) if it satisfies (10).
  2. (ii)
    Let $\varepsilon>0$ be given. Inspired by perez2019subdifferential , we say that $x^{*}$ is an $\varepsilon$-optimization-stationary point of problem (1) if
    $\mathrm{dist}(0,\partial_{\varepsilon}(\Phi+h)(x^{*})+\partial\delta_{\mathcal{M}}(x^{*}))<\varepsilon,$
    with
    $\partial_{\varepsilon}(\Phi+h)(x^{*}):=\{z\in\mathbb{E}\mid\langle z,v\rangle-\varepsilon\leq(\Phi+h)^{\circ}(x^{*};v),~{}\forall v\in\mathbb{E}\}.$
Now we consider defining the game stationarity of problem (1). It is routine to prove that any (local) saddle point $(x^{*},y^{*})$ of problem (1) satisfies
$0\in\nabla_{x}f(x^{*},y^{*})+\partial h(x^{*})+\partial\delta_{\mathcal{M}}(x^{*}),$ (11a)
$0\in\nabla_{y}f(x^{*},y^{*})-\partial g(y^{*})-\partial\delta_{S}(y^{*}),$ (11b)
which yields the following definition of a game-stationary point of problem (1).
Definition 3 (Game-stationary point).
We say that $(x^{*},y^{*})$ is a game-stationary point of problem (1) if it satisfies (11a) and (11b).
Next, we study the relationships between optimization-stationary points and game-stationary points of problem (1).
Proposition 3.
If $x^{*}$ is an optimization-stationary point of the problem (1), then there exists $y^{*}\in S$ such that $(x^{*},y^{*})$ is a game-stationary point. Conversely, if $(x^{*},y^{*})$ is a game-stationary point of problem (1) , then $x^{*}$ is an optimization-stationary point.
Proof.
Suppose that $x^{*}$ is an optimization-stationary point of problem (1). By Proposition (2), $\Phi(x)+h(x)+L_{x}\|x-x^{*}\|^{2}$ is strongly convex. As $0\in\partial\delta_{\mathcal{M}}(x^{*})+\partial\Phi(x^{*})+\partial h(x^{*})$ and $\partial\delta_{\mathcal{M}}(x^{*})=N_{x^{*}}{\mathcal{M}}=\partial\delta_{T_{x^{*}}{\mathcal{M}}}(0)$, then $x^{*}$ is the unique minimum point of the problem
$\min_{x\in A_{x^{*}}\mathcal{M}}\Phi(x)+h(x)+L_{x}\|x-x^{*}\|^{2}$
with $A_{x^{*}}\mathcal{M}:=\{x\mid x=x^{*}+v,~{}v\in T_{x^{*}}{\mathcal{M}}\}$. Therefore, by Sion minimax theorem sionGeneralMinimaxTheorems1958 , there exists $y^{*}\in S$ such that
$f(x^{*},y^{*})+h(x^{*})-g(y^{*})$ $=\inf_{\begin{subarray}{c}x\in A_{x^{*}}{\mathcal{M}},\\ \|x-x^{*}\|\leq 1\end{subarray}}\sup_{y\in S}f(x,y)+h(x)-g(y)+L_{x}\|x-x^{*}\|^{2}$
$=\sup_{y\in S}\inf_{\begin{subarray}{c}x\in A_{x^{*}}{\mathcal{M}},\\ \|x-x^{*}\|\leq 1\end{subarray}}f(x,y)+h(x)-g(y)+L_{x}\|x-x^{*}\|^{2},$
which implies
$0\in\nabla_{x}f(x^{*},y^{*})+\partial h(x^{*})+\partial\delta_{\mathcal{M}}(x^{*}),$
$0\in\nabla_{y}f(x^{*},y^{*})-\partial g(y^{*})-\partial\delta_{S}(y^{*}).$
Conversely, suppose that $(x^{*},y^{*})$ is a game stationary point of problem (1). Then, according to Proposition 2, $y^{*}=\arg\max_{y\in S}f(x^{*},y)-g(y)$ and $\nabla_{x}f(x^{*},y^{*})\in\partial\Phi(x^{*})$. Thus, $0\in\partial\Phi(x^{*})+\partial h(x^{*})+\partial\delta_{\mathcal{M}}(x^{*})$. ∎
Next, we introduce the game stationarity measure. For $(x,y)\in\mathbb{E}_{1}\times\mathbb{E}_{2}$, the game-stationarity measure is defined as
$\mathcal{G}^{\beta}(x,y):=\max\{\|\beta u\|,~{}\mathrm{dist}(0,\nabla_{y}f(x,y)-\partial g(y)-\partial\delta_{S}(y))\},$ (12)
where $\beta>0$ and $u=\arg\min_{v\in T_{x}\mathcal{M}}\langle\nabla_{x}f(x,y),v\rangle+h(x+v)+\frac{\beta}{2}\|v\|^{2}$. This definition is motivated in part by the stationarity measure used in Euclidean minimax problem luHybridBlockSuccessive2020 . It is straightforward to verify that $(x^{*},y^{*})$ is a game-stationary point of problem (1) if and only if there exists $\beta>0$ such that $\mathcal{G}^{\beta}(x^{*},y^{*})=0$. Using this measure, we can also introduce the notion of $\varepsilon$-game-stationary point.
Definition 4.
Let $\varepsilon>0$ be given. We say $(x^{*},y^{*})$ is an $\varepsilon$-game-stationary point of problem (1) if there exists $\beta>0$ such that $\mathcal{G}^{\beta}(x^{*},y^{*})\leq\varepsilon$.
Finally, we examine the connections between $\varepsilon$-optimization-stationary points and $\varepsilon$-game-stationary points of problem (1).
Proposition 4.
Let $\varepsilon>0$ be given. Suppose that $(x^{*},y^{*})$ is an $\varepsilon$-game-stationary point of problem (1) with $\mathcal{G}^{\beta}(x^{*},y^{*})\leq\varepsilon$. Then $x^{*}$ is a $(C\varepsilon)$-optimization-stationary poin, where $C=(2\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}/\beta$ and $\sigma_{y}:=\sup_{y\in S}\left\|y\right\|$.
Proof.
Suppose the point $(x^{*},y^{*})$ is an $\varepsilon$-game-stationary point of problem (1), then there exists $z$ such that $\|z\|=1$ and $\varepsilon z\in\nabla_{y}f(x^{*},y^{*})-\partial g(y^{*})-\partial\delta_{S}(y^{*})$. For $u=\arg\min_{v\in T_{x}\mathcal{M}}\langle\nabla_{x}f(x,y),v\rangle+h(x+v)+\frac{\beta}{2}\|v\|^{2}$, let
$\Phi_{\varepsilon}^{u}(x):=$ $\max_{y\in S}f(x+u,y)-g(y)-\frac{\varepsilon}{2}\|y-(y^{*}-z)\|^{2}$
$+\langle\nabla_{y}f(x^{*},y^{*})-\nabla_{y}f(x^{*}+u,y^{*}),y\rangle.$
By Proposition 2 and the choice of $z$, we have $\nabla\Phi_{\varepsilon}^{u}(x^{*})=\nabla_{x}f(x^{*}+u,y^{*})$ and
$\Phi_{\varepsilon}^{u}(x^{*})+\langle\nabla_{x}f(x^{*}+u,y),x-(x^{*}+u)\rangle\leq\Phi_{\varepsilon}^{u}(x-u)+\frac{L_{x}}{2}\|x-(x^{*}+u)\|^{2}.$
Because
$\Phi(x^{*}+u)-\varepsilon(2\sigma_{y}^{2}+1)-L_{y}\sigma_{y}\|u\|\leq\Phi_{\varepsilon}^{u}(x^{*})$
and
$\Phi_{\varepsilon}^{u}(x-u)\leq\Phi(x)+L_{y}\sigma_{y}\|u\|,$
we arrive at
$\Phi(x^{*}+u)+\langle\nabla_{x}f(x^{*}+u,y),x-(x^{*}+u)\rangle-\varepsilon(2\sigma_{y}^{2}+1)-2L_{y}\sigma_{y}\|u\|$
$\leq$ $\Phi(x)+\frac{L_{x}}{2}\|x-(x^{*}+u)\|^{2}.$
Since $h$ is convex and $L_{h}$-Lipschitz continuous, then
$h(x^{*}+u)+\langle\xi,x-(x^{*}+u)\rangle\leq h(x)$
for all $\xi\in\partial h(x^{*}+u)$. Therefore, by combing the above two inequalities, we obtain
$\nabla_{x}f(x^{*}+u,y^{*})+\partial h(x^{*}+u)\subset\partial_{\hat{\varepsilon}}(\Phi+h)(x^{*}+u),$
with $\hat{\varepsilon}=\varepsilon(2\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}\|u\|$. Furthermore, as $\|\mathcal{G}_{x}^{\beta}(x^{*},y^{*})\|\leq\varepsilon$, we can conclude that $x^{*}$ is a $(C\varepsilon)$-optimization-stationary point with $C=(2\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}/\beta$. ∎
4 The proposed MPGDA algorithm
In this section, we propose our MPGDA method for solving problem (1) and establish its iteration complexity. Our algorithm development is partially motivated by heApproximationProximalGradient2024 ; luHybridBlockSuccessive2020 ; xuRiemannianAlternatingDescent2024 . For the $k$-th iteration, we introduce the following value function:
$Q_{k}(x):=h(x)+\Phi_{k}(x)$ (13)
where $\Phi_{k}:\mathbb{E}_{1}\to\mathbb{R}$ is defined at $x\in\mathbb{E}_{1}$ by
$\Phi_{k}(x):=\max_{y\in S}\{f(x,y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}\}.$ (14)
Here, $y_{k}$ denotes the iterate with respect to $y$ obtained in the previous iteration, $\gamma_{k}>0$ is the regularization parameter, and $\rho_{k}>0$ is the proximal parameter. Moreover, we choose $\gamma_{k}=\gamma_{0}/k^{1/3}$ with some $\gamma_{0}>0$ and assume the non-increasing sequence $\{\rho_{k}\}$ to be summable, i.e.,
$\{\rho_{k}\}\in\mathcal{S}:=\{\{\rho_{k}\}\mid\sum_{k=0}^{\infty}\rho_{k}<+\infty,~{}~{}\rho_{k}\geq\rho_{k+1}>0,k=0,1,\dots~{}\}.$
We further introduced the mapping $\bar{y}_{k}:\mathbb{E}_{1}\to\mathbb{E}_{2}$ defined by
$\bar{y}_{k}(x):=\arg\max_{y\in S}\{f(x,y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}\}.$ (15)
Then it holds that $Q_{k}(x)=F(x,\bar{y}_{k}(x))-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x)\|^{2}-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x)-y_{k}\|^{2}$ and in view of Danskin’s theorem (danskinTheoryMaxMinApplications1966, , Theorem D1), $\Phi_{k}$ is smooth on $\mathbb{E}_{1}$ and $\nabla\Phi_{k}(x)=\nabla_{x}f(x,\bar{y}_{k}(x))$.
Now, at the $k$-th iteration, the proposed MPGDA method computes the new iterate by approximately solving the subproblem:
$\min_{x\in\mathcal{M}}Q_{k}(x).$ (16)
Inspired by the structure of $Q_{k}$ and the ManPG algorithm (chenProximalGradientMethod2020, , Algorithm 4.1), we perform Riemannian manifold proximal gradient descent steps for $T_{k}$ times to generate $x_{k+1}$ with $T_{k}$ being an integer in $[1,\overline{T}]$ where $\overline{T}$ is a predetermined positive integer. Then $y_{k+1}$ is updated as $\bar{y}_{k}(x_{k+1})$ with $\bar{y}_{k}$ defined in (15). Notice that the update of $y$ can also be interpreted as performing a proximal ascent step on the function $f(x_{k+1},y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}$. We formally presented the proposed method in Algorithm 1.
Algorithm 1 Manifold Proximal Gradient Descent Ascent (MPGDA) for problem (1)
0:  $c_{1}>0$, $\eta\in(0,1)$, $\gamma_{0}>0$, $x_{0}\in\mathcal{M}$, $y_{0}\in\mathrm{dom}(g)\cap S$, $0<l_{\min}<l_{\max}$, $\{\rho_{k}\}\in\mathcal{S}$, $\{\gamma_{k}=\gamma_{0}/k^{1/3}:k\in N\}$, $\{T_{k}\}\subset[1,\overline{T}]$;
1:  for $k=0,\dots$ do
2:     Set $x_{k,0}=x_{k}$,
3:     for $i=0,\dots,T_{k}-1$ do
4:        Choose $l_{k,i}\in[l_{\min},l_{\max}]$ and set $\beta_{k,i}=l_{k,i}/(\rho_{k}+\gamma_{k})$.
5:        Compute
$v_{k,i}:=\arg\min_{v\in T_{x_{k,i}}\mathcal{M}}\langle\nabla_{x}f(x_{k,i},\bar{y}_{k}(x_{k,i})),v\rangle+h(x_{k,i}+v)+\frac{\beta_{k,i}}{2}\left\|v\right\|^{2},$ (17)
6:        Find the smallest non-negative integer $j$ such that
$Q_{k}(R_{x_{k,i}}(\eta^{j}v_{k,i}))\leq Q_{k}(x_{k,i})-c_{1}\eta^{j}\beta_{k,i}\|v_{k,i}\|^{2}+2\rho_{k}\sigma_{y}^{2}$ (18)
and update $x_{k,i+1}=R_{x_{k,i}}(\eta^{j}v_{k,i})$
7:     end for
8:     $x_{k+1}=x_{k,T_{k}},~{}y_{k+1}=\bar{y}_{k}(x_{k+1})$.
9:  end for
10:  return $(x_{k+1},y_{k+1})$.
Before proceeding, we make two remarks on the iteration procedure of MPGDA algorithm. First, given $x_{k,i}\in\mathcal{M}$, MPGDA algorithm computes a descent direction $v_{k,i}$ of $Q_{k}$ restricted to the tangent space $T_{x_{k,i}}\mathcal{M}$ via tackling the linear constrained strongly convex optimization problem (17), which can usually be efficiently solved, see (li2024proximal, , Section 4), for example. Second, in practice, we make some choices of parameters $l_{k,i}$ and $\rho_{k}$, which enhance the efficiency of MPGDA algorithm in our test and also satisfy $l_{k,i}\in[l_{\min},l_{\max}]$ and $\rho_{k}\in S$. For parameters $l_{k,i}$, we adopt the Riemannian Barzilai-Borwein (BB) stepsize iannazzo2018riemannian ; 2013AWen as follows
$l_{k,i}=\begin{cases}\min\{\max\{l_{\min},(\rho_{k}+\gamma_{k})|\frac{\langle\Delta X,\Delta G\rangle}{\|\Delta X\|^{2}}|)\},l_{\max}\},&if~{}\Delta X\neq 0,\\ l_{\max},&if~{}\Delta X=0,\end{cases}$
where $\Delta X=X_{k,i}-X_{k,i-1}$ and $\Delta G=\mathrm{grad}\Phi_{k}(x_{k,i})-\mathrm{grad}\Phi_{k}(x_{k,i-1})$. Moreover, inspired by xuRiemannianAlternatingDescent2024 , for given constant $\theta>1$, $\tau_{1}\in(0,1)$ and $\tau_{2}\in(0,1)$, we set $\rho_{0}=\xi_{0}$ and for $k\geq 1$,
$\rho_{k}=\frac{\xi_{k}}{k^{\theta}}~{}~{}\text{with}~{}~{}\xi_{k}=\begin{cases}\tau_{2}\xi_{k-1},&\text{if}~{}~{}\delta_{k}\geq\tau_{1}\delta_{k-1},\\ \xi_{k-1},&\text{else},\end{cases}$ (19)
where $\delta_{k}$ is defined as
$\delta_{k}=\|\gamma_{k-1}y_{k}+\rho_{k-1}(y_{k}-y_{k-1})\|_{\infty}.$
In what follows, we conduct convergence analysis for MPGDA algorithm. First, we present a technical lemma that guarantees the boundedness of the mapping $\bar{y}_{k}$.
Lemma 1.
Let the mapping $\bar{y}_{k}:\mathbb{E}_{1}\to\mathbb{E}_{2}$ be defined in (15) for $k\in N$. Then for all $x$, $\bar{x}\in\mathbb{E}_{1}$ and $y\in\mathbb{E}_{2}$ , it holds that
$f(x,y)-g(y)$
$\leq f(x,\bar{y}_{k}(\bar{x}))-g(\bar{y}_{k}(\bar{x}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(\bar{x})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(\bar{x})\|^{2}+\frac{\gamma_{k}}{2}\|y\|^{2}$
$+\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}+\frac{L_{y}^{2}}{2(\rho_{k}+\gamma_{k})}\|x-\bar{x}\|^{2}.$
Proof.
According to the definition of $\bar{y}_{k}(\bar{x})$, we have
$0\in\nabla_{y}f(\bar{x},\bar{y}_{k}(\bar{x}))-\partial g(\bar{y}_{k}(\bar{x}))-\partial\delta_{S}(\bar{y}_{k}(\bar{x}))-\rho_{k}(\bar{y}_{k}(\bar{x})-y_{k})-\gamma_{k}\bar{y}_{k}(\bar{x}).$
By subgradient inequality, we obtain
$f(x,y)-g(y)$
$\leq f(x,\bar{y}_{k}(\bar{x}))-g(\bar{y}_{k}(\bar{x}))+\langle\nabla_{y}f(x,\bar{y}_{k}(\bar{x}))-\nabla_{y}f(\bar{x},\bar{y}_{k}(\bar{x})),y-\bar{y}_{k}(\bar{x})\rangle$
$+\langle\rho_{k}(\bar{y}_{k}(\bar{x})-y_{k})+\gamma_{k}\bar{y}_{k}(\bar{x}),y-\bar{y}_{k}(\bar{x})\rangle.$
Since $2\langle a,a-b\rangle=\left\|a\right\|^{2}+\left\|a-b\right\|^{2}-\left\|b\right\|^{2}$,by Cauchy-Schwarz inequality, we arrive at
$f(x,y)-g(y)$
$\leq f(x,\bar{y}_{k}(\bar{x}))-g(\bar{y}_{k}(\bar{x}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(\bar{x})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(\bar{x})\|^{2}+\frac{\gamma_{k}}{2}\|y\|^{2}$
$+\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}+\frac{L_{y}^{2}}{2(\rho_{k}+\gamma_{k})}\|x-\bar{x}\|^{2}.$
Next, we prove that the descent condition (18) must be satisfied after a unified finite number of iterations for all $k\in N$ and $i\in[0,T_{k}-1]$, which ensures that the proposed Algorithm 1 is well-defined.
Proposition 5.
Consider Algorithm 1. There exists a constant $\bar{\alpha}>0$ such that for all $k\in N$ and $0<\alpha\leq\min\{1,~{}\bar{\alpha}\}$, it holds that
$Q_{k}(R_{x_{k,i}}(\alpha v_{k,i}))\leq Q_{k}(x_{k,i})-\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}+2\rho_{k}\sigma_{y}^{2}$
for all $i=0,1,\dots,T_{k}-1$.
Proof.
By (chenProximalGradientMethod2020, , Lemma 5.2), for any $0<\alpha\leq\min\{1,~{}\frac{\beta_{k,i}}{2M_{2}(L_{h}+G)+M_{1}^{2}L_{x}}\}$,
$f(x_{\alpha},\bar{y}_{k}(x_{k,i}))+h(x_{\alpha})\leq f(x_{k,i},\bar{y}_{k}(x_{k,i}))+h(x_{k,i})-\frac{\alpha\beta_{k,i}}{2}\|v_{k,i}\|^{2},$
where $x_{\alpha}=R_{x_{k,i}}(\alpha v_{k,i})$ and $G=\sup_{x\in\mathcal{M},y\in\mathrm{dom}(g)\cap S}\|\nabla_{x}F(x,y)\|$.
According to Lemma 1, we have
$f(x_{\alpha},\bar{y}_{k}(x_{\alpha}))-g(\bar{y}_{k}(x_{\alpha}))$
$\leq f(x_{\alpha},\bar{y}_{k}(x_{k,i}))-g(\bar{y}_{k}(x_{k,i}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{k,i})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{k,i})\|^{2}+\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{\alpha})\|^{2}$
$+\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{\alpha})-y_{k}\|^{2}+\frac{L_{y}^{2}}{2(\rho_{k}+\gamma_{k})}\|x_{\alpha}-x_{k,i}\|^{2}.$
By summing these inequalities with the inequality
$\|x_{\alpha}-x_{k,i}\|\leq\alpha M_{1}\|v_{k,i}\|$
in Proposition 1, we derive the following inequality:
$F(x_{\alpha},\bar{y}_{k}(x_{\alpha}))-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{\alpha})\|^{2}$
$\leq F(x_{k,i},\bar{y}_{k}(x_{k,i}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{k,i})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{k,i})\|^{2}$
$+\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{\alpha})-y_{k}\|^{2}-(\frac{\alpha\beta_{k,i}}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2(\rho_{k}+\gamma_{k})})\|v_{k,i}\|^{2}.$
Since $\beta_{k,i}=l_{k,i}/(\rho_{k}+\gamma_{k})$, for any $0<\alpha\leq\frac{l_{k,i}}{2M_{1}^{2}L_{y}^{2}}$, we have
$(\frac{\alpha\beta_{k,i}}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2(\rho_{k}+\gamma_{k})})\|v_{k,i}\|^{2}$ $=\beta_{k,i}(\frac{\alpha}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2l_{k,i}})\|v_{k,i}\|^{2}\geq\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}.$
Note that $Q_{k}(x)=F(x,\bar{y}_{k}(x))-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x)\|^{2}-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x)-y_{k}\|^{2}$, then for any $0<\alpha\leq\min\{1,\frac{l_{\min}}{(\rho_{0}+\gamma_{0})(2M_{2}(L_{h}+G)+M_{1}^{2}L_{x})},\frac{l_{\min}}{2M_{1}^{2}L_{y}^{2}}\}$, we have
$Q_{k}(x_{\alpha})\leq Q_{k}(x_{k,i})-\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}+2\rho_{k}\sigma_{y}^{2}$
for all $k\geq 0$. ∎
We are now ready to establish the iteration complexity for MPGDA algorithm. Let $\varepsilon>0$ be a given target accuracy, we define
$T(\varepsilon):=\min\{k\mid\mathcal{G}^{\beta_{k,0}}(x_{k,0},\bar{y}_{k}(x_{k,0}))<\varepsilon\}$
where $\mathcal{G}^{\beta}(\cdot,\cdot)$ is defined in (12). In other words, $T(\varepsilon)$ stands for the minimal number of outer iterations needed for MPGDA algorithm to obtain a $\varepsilon$-game-stationary point. We further define two constants $\overline{Q}$ and $\underline{Q}$ as
$\overline{Q}$ $:=\sup\{{F(x,y)\mid x\in\mathcal{M},y\in\mathrm{dom}(g)\cap S}\},$
$\underline{Q}$ $:=\inf\{F(x,y)-\frac{\rho_{0}}{2}\|y-y_{0}\|^{2}-\frac{\gamma_{0}}{2}\|y\|^{2}\mid x\in\mathcal{M},y\in\mathrm{dom}(g)\cap S\}.$
Theorem 4.1.
Given $\varepsilon>0$, let $\{(x_{k},y_{k})\}$ be generated by Algorithm 1. Then, there exists a positive $\bar{\alpha}>0$ such that the following inequality holds:
$T(\varepsilon)\leq\frac{\max\{(2\rho_{0}+\gamma_{0})\sigma_{y},C\}}{\varepsilon^{3}}$
with
$C=\frac{\sqrt{2(\overline{Q}-\underline{Q}+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2})l_{\max}}}{\sqrt{3\bar{\alpha}\eta\gamma_{0}}}.$
Here, all the parameters $\rho_{0},~{}\gamma_{0},~{}\sigma_{y},~{}\overline{Q},~{}\underline{Q},~{}\overline{T},~{}S_{\rho},~{}l_{\max},~{}\eta$ are defined as in the relevant context of the paper.
Proof.
Let $\beta_{k}:=\beta_{k,0},~{}v_{k}:=v_{k,0}$. Since $\gamma_{k+1}\leq\gamma_{k}$ and $\rho_{k+1}\leq\rho_{k}$, by the definition of $Q_{k}$, we have
$Q_{k+1}(x_{k+1})$ $=h(x_{k+1})+\max_{y\in S}f(x_{k+1},y)-g(y)-\frac{\gamma_{k+1}}{2}\|y\|^{2}-\frac{\rho_{k+1}}{2}\|y-y_{k+1}\|^{2}$
$\leq h(x_{k,T_{k}})+\max_{y\in S}f(x_{k,T_{k}},y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}+4\rho_{k}\sigma_{y}^{2}$
$=Q_{k}(x_{k,T_{k}})+4\rho_{k}\sigma_{y}^{2}.$
Furthermore, according to Proposition 5, there exists $\bar{\alpha}>0$ such that
$Q_{k+1}(x_{k+1})\leq Q_{k}(x_{k})+2(T_{k}+2)\rho_{k}\sigma_{y}^{2}-\bar{\alpha}\eta\beta_{k}\|v_{k}\|^{2}.$
Summing the above inequality over $k=K_{1},K_{1}+1,\dots,K$ yields
$\sum_{k=K_{1}}^{K}\bar{\alpha}\eta\beta_{k}\|v_{k}\|^{2}\leq{Q}_{K_{1}}(x_{K_{1}})-{Q}_{K}(x_{K})+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2}$ (20)
with $S_{\rho}=\sum_{k=0}^{\infty}\rho_{k}<\infty$.
Since $\beta_{k}\leq\frac{l_{\max}}{\gamma_{0}k^{-1/3}}$, then
$\sum_{k=K_{1}}^{K}\bar{\alpha}\eta\beta_{k}\|v_{k}\|^{2}\geq(\frac{\bar{\alpha}\eta\gamma_{0}}{l_{\max}}\sum_{k=K_{1}}^{K}k^{1/3})\min_{K_{1}\leq k\leq K}\|\beta_{k}v_{k}\|^{2}.$ (21)
Note that
[…]
let $K=\lceil{2^{3/2}K_{1}}\rceil$ , then we obtain
$\sum_{K_{1}\leq k\leq\lceil{2^{3/2}K_{1}}\rceil}\frac{1}{k^{1/3}}\geq\frac{3}{2}K_{1}^{2/3}.$ (22)
Therefore, combining (20), (21) and (22), with $Q_{K}(x_{K_{1}})\leq\overline{Q}$ and $Q_{\lceil{2^{3/2}K_{1}}\rceil}(x_{\lceil{2^{3/2}K_{1}}\rceil})\geq\underline{Q}$, we have
$\min_{K_{1}\leq k\leq\lceil{2^{3/2}K_{1}}\rceil}\|\beta_{k}v_{k}\|^{2}\leq\frac{2l_{\max}}{3\bar{\alpha}\eta\gamma_{0}K_{1}^{2/3}}(\overline{Q}-\underline{Q}+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2}).$ (23)
In view of definition (15) and the associated first-order optimality condition, for $\bar{y}_{k}(x_{k,0})$, we have that
$\rho_{k}(\bar{y}_{k}(x_{k,0})-y_{k})+\gamma_{k}\bar{y}_{k}(x_{k,0})\in\nabla_{y}f(x_{k,0},\bar{y}_{k}(x_{k,0}))-\partial g(\bar{y}_{k}(x_{k,0}))-\partial\delta_{S}(\bar{y}_{k}(x_{k,0})).$
Invoking this inclusion, we further obtain that
$\mathrm{dist}(0,\nabla_{y}f(x_{k,0},\bar{y}_{k}(x_{k,0}))-\partial g(\bar{y}_{k}(x_{k,0}))-\partial\delta_{S}(\bar{y}_{k}(x_{k,0})))$ (24a)
$\leq$ $\|\rho_{k}(\bar{y}_{k}(x_{k,0})-y_{k})+\gamma_{k}\bar{y}_{k}(x_{k,0})\|$ (24b)
$\leq$ $\sigma_{y}(2\rho_{k}+\gamma_{k}).$ (24c)
Finally, by the definition of $\rho_{k}$, $\gamma_{k}$ and $\mathcal{G}^{\beta_{k}}$ , combining (23) and (24), we conclude that
$\min_{K_{1}\leq k\leq\lceil 2^{3/2}K_{1}\rceil}\mathcal{G}^{\beta_{k}}(x_{k,0},\bar{y}_{k}(x_{k,0}))\leq\frac{\max\{(2\rho_{0}+\gamma_{0})\sigma_{y},C\}}{K_{1}^{1/3}}$
with
$C=\frac{\sqrt{2(\overline{Q}-\underline{Q}+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2})l_{\max}}}{\sqrt{3\bar{\alpha}\eta\gamma_{0}}},$
and the proof is finished. ∎
Theorem 4.1 demonstrates that the proposed MPGDA algorithm can find an $\varepsilon$-game-stationary point for problem (1) within $O(\varepsilon^{-3})$ iterations. In view of Proposition 4, this also indicates that the iteration complexity of the MPGDA algorithm for returning an $\varepsilon$-optimization-stationary point is $\mathcal{O}(\varepsilon^{-3})$.
5 Numerical experiments
In this section, we demonstrate the efficiency of the proposed MPGDA algorithm by applying it to FSPCA problem (4) and SSC problem (7). All the experiments are implemented in MATLAB R2019a and conducted on a standard PC with 3.40GHz Intel(R) Core(TM) i5-7500 CPU and 16GB of RAM.
We first specify some implementation details of the proposed algorithm. Throughout the tests, we set $c_{1}=10^{-4}$, $\eta=0.1$, $l_{\min}=10^{-16}$, $l_{\max}=10^{16}$, $\tau_{1}=0.999$ and $\tau_{2}=0.9$. The parameters $\gamma_{0}$, $\xi_{0}>0$, $\theta>1$ and $\{T_{k}\}$ vary across the two application problems. Since $h$ is the $\ell_{1}$-norm in both problems, the semismooth Newton method developed in (chenProximalGradientMethod2020, , Section 4.2) is used to solve subproblem (17). We utilize the QR factorization as the retraction on the Stiefel manifold 2008Optimization .
5.1 Experiments on FSPCA
In this subsection, we compare the proposed MPGDA algorithm with the constraint relaxed majorization-minimization (CRMM) algorithm (babuFairPrincipalComponent2023, , Algorithm 3) for solving FSPCA problem (4). The CRMM algorithm first relaxes the semi-orthogonality constraint $X^{T}X=I$ into a linear matrix inequality constraint and then applies some majorization-minimization iteration procedure, which requires solving a quadratic program at each iteration.
We perform tests on synthetic datasets and a real dataset where the samples belong to two groups. The synthetic datasets are generated following similar settings to (2016SparseGeneralizedEigenvalueProblem, , Section 5.1). In detail, for the $i$-th group, we randomly generate $m_{i}=200$ samples with $d=40$ features according to Gaussian distribution with mean $\mu^{i}\in\mathbb{R}^{d}$ and covariance $\Sigma\in\mathbb{R}^{d\times d}$. We set $\mu_{j}^{1}=0,\ \mu_{j}^{2}=1/3$, for $j\in\{2,4,\cdots,200\}$, and $\mu_{j}^{i}=0$ otherwise. Meanwhile, the covariance matrix $\Sigma$ is chosen as a block diagonal matrix with five $8\times 8$ blocks, where the $(j,j^{\prime})$-th entry of each block is $0.8^{|j-j^{\prime}|}$. For the real-world data, we use the default credit dataset yeh2009creditdata and preprocess it as done in samadiPriceFairPCA . We set the regularization parameter $\mu=0.1$ in our tests. Furthermore, the CRMM algorithm is terminated if the relative error between the solutions obtained over two successive iterations is less than $10^{-6}$. For MPGDA algorithm, we set $\gamma_{0}=10^{-6}$, $\xi_{0}=4\sqrt{r}*10^{4}$, $\theta=1.5$, $T_{k}\equiv 15$, and terminate the algorithm if it satisfies:
$\mathcal{G}^{\beta_{k,0}}(x_{k},y_{k})<\varepsilon$ (25)
with $\varepsilon=10^{-6}$. Both compared algorithms start from the same random initial points on $\mathrm{St}(d,r)$ and the maximum iteration unmber of them are set as 1000.
We present the average results over 50 runs on different synthetic datasets and results for the default credit dataset in Table 1. We report the objective values (Obj), and CPU time in seconds (Time) and the number of iterations (Iter). It is observed that MPGDA algorithm achieves similar objective values to the CRMM algorithm while requiring much less computational time. Specially, our proposed algorithm is on average more than 20 times faster than the CRMM algorithm.
Table 1: Results for FSPCA
CRMM MPGDA
Dataset r Obj. Time (s) Iter. Obj. Time (s) Iter.
Synthetic 2 $-9.795$ $0.269$ $201$9.808 0.007 167
3 $-14.380$ $0.344$ 155 -14.382 0.010 $192$
4 $-18.714$ $0.443$ 162 -18.745 0.012 $217$
5 $-22.774$ $0.209$ 83 -22.809 0.008 $197$
Credit 2 $-9.531$ $4.014$ 93 -9.742 0.133 $120$
3 $-10.557$ $15.692$ $271$10.802 0.104 91
4 $-11.637$ $8.977$ $142$11.923 0.010 5
5 $-12.572$ $20.058$ $277$12.724 0.012 4
5.2 Experiments on SSC
In this subsection, we compare the proposed MPGDA method with the RADA-RGD and RADA-PGD algorithms in xuRiemannianAlternatingDescent2024 for solving the SSC problem. Note that in xuRiemannianAlternatingDescent2024 , the original SSC problem (5) is reformulated into the following smooth minimax problem on the Grassmann manifold so that the RADA-PGD and RADA-RGD algorithms can be directly applied:
$\min_{Z\in\mathrm{Gr}(N,p)}\max_{\|Y\|_{\infty}\leq\mu}\{\langle L,Z\rangle+\langle Y,Z\rangle\}.$ (26)
Here, $\mathrm{Gr}(N,p)=\{XX^{\top}\in\mathbb{R}^{N\times N}\mid X\in\mathrm{St}(N,p)\}$ denotes the Grassmann manifold AGrassmannmanifoldhandbook2024A . The codes of these two algorithms for solving (26) are downloaded from: https://github.com/XuMeng00124/RADAopt.
In the tests, we generate synthetic datasets as in 2016ConvexSparseSpectralClustering , where the data points $\{a_{i}\}_{i=1}^{N}$ are independently drawn from the standard Gaussian distribution and $W_{ij}=|\langle a_{i},a_{j}\rangle|$. For the two RADA algorithms, we adopt the parameters as suggested in (xuRiemannianAlternatingDescent2024, , Section 6.3). Specifically, they are terminated once an $\varepsilon$-game-stationary point of problem (26) is found with $\varepsilon=10^{-4}$ and the initial point is set as $X_{0}X_{0}^{\top}$, where $X_{0}$ consists of $p$ eigenvectors associated with the $p$ smallest eigenvalues of $L$. For MPGDA algorithm, we set $\gamma_{0}=10^{-5}$, $\xi_{0}=\sqrt{p}N^{2}$, $\theta=2$ and $T_{k}\equiv 3$. Moreover, we choose the aforementioned $X_{0}$ as the initial point of MPGDA algorithm and terminate the method if it satisfies (25) with $\varepsilon=10^{-4}$. The maximum number of iterations for all the compared algorithms is set as 1000.
Table 2: Results for SSC
Config Para $(p/\mu)$ RADA-PGD RADA-RGD MPGDA
Obj. Time Iter. Obj. Time Iter. Obj. Time Iter.
$\mu=0.1$ $N=200$ 2 2.187 $0.513$ $265$ 2.187 $0.796$ $153$ 2.187 0.114 65
4 $4.374$ $0.530$ $259$ 4.373 $0.812$ $157$ $4.374$ 0.116 64
6 6.560 $0.521$ $224$ 6.560 $0.855$ $164$ 6.560 0.117 63
8 8.747 $0.580$ $198$ 8.747 $0.861$ $166$ 8.747 0.129 64
10 10.934 $0.621$ $180$ 10.934 $0.995$ $167$ 10.934 0.129 65
$p=5$ $N=200$ 0.1 5.467 $0.535$ $244$ 5.467 $0.868$ $164$ 5.467 0.122 63
0.2 5.967 $0.556$ $263$ 5.967 $0.934$ $178$ 5.967 0.133 71
0.5 7.467 $0.593$ $286$ 7.467 $1.071$ $202$ 7.467 0.174 86
1.0 9.967 $0.616$ $300$ 9.967 $1.188$ $228$ 9.967 0.205 95
The computational results averaged over 50 synthetic data sets with $N=200$ data points are summarized in Table 2, where we report the obtained objective function value of problem (5) (Obj.), the CPU time in seconds (Time), and number of outer iterations (Iter.). We observe that all the compared algorithms attain comparable objective function value for different number of groups and regularization parameters $\mu$. In addition, the proposed MPGDA method consistently outperform the two RADA algorithms in terms of CPU time, which again validates its computational efficiency.

### v2

@@
 Owing to the nonsmoothness of the objective function, existing minimax manifold optimization methods cannot be directly applied to solve this problem.
-We propose a manifold proximal gradient descent ascent (MPGDA) algorithm for solving the problem.
-At each iteration, the proposed algorithm alternately performs one or multiple manifold proximal gradient descent steps and a proximal ascent step.
-We prove that the MPGDA algorithm can find an $\varepsilon$-game-stationary point and an $\varepsilon$-optimization-stationary point of the considered problem within $\mathcal{O}(\varepsilon^{-3})$ iterations.
-Numerical experiments on fair sparse PCA and sparse spectral clustering are conducted to demonstrate the advantages of the MPGDA algorithm.
+We propose two manifold proximal gradient descent ascent (MPGDA) algorithms for solving the problem.
+The first algorithm alternatively performs one or multiple manifold proximal gradient descent steps and a proximal ascent step at each iteration, and we prove that it can find an $\varepsilon$-game-stationary point and an $\varepsilon$-optimization-stationary point within $\mathcal{O}(\varepsilon^{-3})$ outer iterations.
+The second algorithm alternatively performs one manifold proximal gradient descent step and a proximal gradient ascent step, and we show that it can reach an $\varepsilon$-game-stationary point and an $\varepsilon$-optimization-stationary point within $\mathcal{O}(\varepsilon^{-4})$ outer iterations.
+Numerical experiments on an analytic example, fair sparse PCA, and sparse spectral clustering are conducted to illustrate the advantages of the proposed algorithms.
 Keywords:
@@
 90C26 90C30 65K05
+journal: Journal of Scientific Computing
 1 Introduction
@@
 $\min_{x\in\mathcal{M}}\max_{y\in S}F(x,y):=f(x,y)+h(x)-g(y),$ (1)
-where $\mathcal{M}$ is a Riemannian manifold embedded in a finite-dimensional Euclidean space $\mathbb{E}_{1}$, and $S$ is a convex set in another finite-dimensioanl Euclidean space $\mathbb{E}_{2}$, $f:\mathbb{E}_{1}\times\mathbb{E}_{2}\to\mathbb{R}$ is a continuously differentiable function and is concave with respect to y, $h:\mathbb{E}_{1}\to\mathbb{R}$ and $g:\mathbb{E}_{2}\to\mathbb{R}$ are some proper closed convex (possibly nonsmooth) functions.
+where $\mathcal{M}$ is a Riemannian manifold embedded in a finite-dimensional Euclidean space $\mathbb{E}_{1}$, and $S$ is a convex set in another finite-dimensional Euclidean space $\mathbb{E}_{2}$, $f:\mathbb{E}_{1}\times\mathbb{E}_{2}\to\mathbb{R}$ is a continuously differentiable function and is concave with respect to y, $h:\mathbb{E}_{1}\to\mathbb{R}$ and $g:\mathbb{E}_{2}\to\mathbb{R}$ are some proper closed convex (possibly nonsmooth) functions.
 Throughout this paper, we adopt the following assumptions for problem (1).
 {assumption}
-$(i)$ The manifold $\mathcal{M}$ is compact.
+$(i)$ The manifold $\mathcal{M}$ and the convex set $S$ are compact. Let $\sigma_{y}:=\sup_{y\in S}\left\|y\right\|<\infty$.
 $(ii)$ The function $f$ has Lipschitz continuous gradient with respect to $x$, i.e., there exists a Lipschitz constant $L_{x}>0$ such that
@@
 $(iv)$ The function $h$ is $L_{h}$ Lipschitz continuous.
+$(v)$ For all $u\in\mathbb{E}_{1},~w\in\mathbb{E}_{2}$ and $\alpha>0$, the following two proximal problems can be solved exactly and efficiently:
+$\min_{x\in\mathbb{E}_{1}}h(x)+\frac{1}{2\alpha}\|x-u\|^{2},$ (2a)
+$\min_{y\in S}g(y)+\frac{1}{2\alpha}\|y-w\|^{2}.$ (2b)
 Problem (1) has many important applications in machine learning and signal processing.
+Note that if $f(x,y)$ is linear with respect to $y$ for fixed $x$, then $f(x,y)=\langle\mathcal{A}(x),y\rangle$ for some smooth operator $\mathcal{A}:\mathbb{E}_{1}\to\mathbb{E}_{2}$.
 Below, we present two representative examples and refer the interested readers to chenProximalGradientMethod2020; BriefIntroductiontoManifoldOptimization for more examples.
 Example 1.
-Fair Sparse Principle Component Analysis (FSPCA).
-The classical PCA Hotelling1933Analysis is one of the most widely used dimensional reduction techniques, which minimizes the total reconstruction error over the whole samples.
+Fair Sparse Principal Component Analysis (FSPCA).
+The classical PCA Hotelling1933Analysis is one of the most widely used dimensionality reduction techniques, which minimizes the total reconstruction error over the whole samples.
 However, in real-life applications, the data samples come from diverse classes and PCA may exhibit a higher reconstruction error for certain classes than the others.
@@
 $\min_{X\in\mathrm{St}(d,r)}\max_{i=1,\dots,n}-\mathrm{Tr}(X^{\top}A_{i}^{\top}A_{i}X),$ (3)
-where $r<d$ is the number of principle components, $\mathrm{Tr}(\cdot)$ represents the trace and $\mathrm{St}(d,r):=\{X\in\mathbb{R}^{d\times r}|X^{\top}X=I_{r}\}$ with $I_{r}$ denoting the $r\times r$ identity matrix.
-Recently, the fair sparse PCA is further proposed to promote sparsity in the principle components babuFairPrincipalComponent2023 as follows:
+where $r<d$ is the number of principal components, $\mathrm{Tr}(\cdot)$ represents the trace and $\mathrm{St}(d,r):=\{X\in\mathbb{R}^{d\times r}|X^{\top}X=I_{r}\}$ with $I_{r}$ denoting the $r\times r$ identity matrix.
+Recently, the fair sparse PCA is further proposed to promote sparsity in the principal components babuFairPrincipalComponent2023 as follows:
 $\min_{X\in\mathrm{St}(d,r)}\max_{i=1,\dots,n}-\mathrm{Tr}(X^{\top}A_{i}^{\top}A_{i}X)+\mu\left\|X\right\|_{1},$ (4)
 where $\mu>0$ is a weighting parameter and $\left\|X\right\|_{1}:=\sum_{i,j}\left|x_{ij}\right|$ is the $\ell_{1}$-norm of the matrix $X$.
-In view of (Beck2017First, , Example 4.10), problem (3) can be equivalently reformulated into
+In view of (Beck2017First, Example 4.10), problem (4) can be equivalently reformulated into
 $\min_{X\in\mathrm{St}(d,r)}\max_{y\in\Delta_{n}}-\sum_{i=1}^{n}y_{i}\mathrm{Tr}(X^{\top}A_{i}^{\top}A_{i}X)+\mu\left\|X\right\|_{1}.$ (5)
@@
 In spectral clustering, a symmetric affinity matrix $W=[W_{ij}]_{N\times N}$ is constructed, where $W_{ij}\geq 0$ quantifies the pairwise similarity between samples $a_{i}$ and $a_{j}$.
-To promote sparsity and interpretability, the sparse spectral clustering framework is introduced in 2016ConvexSparseSpectralClustering , as follows
+To promote sparsity and interpretability, the sparse spectral clustering framework is introduced in 2016ConvexSparseSpectralClustering, as follows:
 $\min_{X\in\mathrm{St}(N,p)}\langle L,XX^{\top}\rangle+\mu\|XX^{\top}\|_{1},$ (6)
@@
 In recent years, nonconvex-concave (NC-C) minimax problems have received a tremendous amount of attention as they have found broad applications in diverse areas.
-Many studies focus on NC-C problems in the setting where both the constraints on $x$ and $y$ are closed convex sets and numerous methods have been developed, see, for example, linGradientDescentAscent ; linOptimalAlgorithmsMinimax ; nouiehedSolvingClassNonConvex ; ostrovskiiEfficientSearchFirstOrder2021 ; panEfficientAlgorithmNonconvexlinear2021a ; 2022Zeroth ; thekumparampilEfficientAlgorithmsSmooth ; xuUnifiedSingleloopAlternating2023 ; yangFasterSingleloopAlgorithms ; zhangSingleLoopSmoothedGradient .
+Many studies focus on NC-C problems in the setting where both the constraints on $x$ and $y$ are closed convex sets and numerous methods have been developed, see, e.g., linGradientDescentAscent; linOptimalAlgorithmsMinimax; nouiehedSolvingClassNonConvex; ostrovskiiEfficientSearchFirstOrder2021; panEfficientAlgorithmNonconvexlinear2021a; 2022Zeroth; thekumparampilEfficientAlgorithmsSmooth; xuUnifiedSingleloopAlternating2023; yangFasterSingleloopAlgorithms; zhangSingleLoopSmoothedGradient.
 In contrast, there are relatively few works studying NC-C minimax problems in the setting where the constraint on $x$ is a Riemannian submanifold and the constraint on $y$ is a closed convex set.
-Among these works, Huang and Gao huangGradientDescentAscent2023 propose a Riemannian gradient descent-ascent algorithm for a smooth nonconvex-strongly concave (NC-SC) minimax problems and establish its iteration complexity of $\mathcal{O}(\varepsilon^{-2})$ for achieving $\varepsilon$-stationarity.
+Among these works, Huang and Gao huangGradientDescentAscent2023 propose a Riemannian gradient descent-ascent algorithm for a smooth nonconvex-strongly concave (NC-SC) minimax problem and establish its iteration complexity of $\mathcal{O}(\varepsilon^{-2})$ for achieving $\varepsilon$-stationarity.
 Later, Xu et al.
 xuEfficientAlternatingRiemannian2024 present an alternating Riemannian/projected gradient descent-ascent (ARPGDA) algorithm with an iteration complexity of $\mathcal{O}(\varepsilon^{-3})$ for finding an $\varepsilon$-stationary point of smooth Riemannian nonconvex-linear (NC-L) minimax problem.
-Very recently, for problem (1) with additional assumptions that $h$ is smooth (but not necessarily convex) and $f(x,y)$ is linear with respect to $y$ for fixed $x$, the work xuRiemannianAlternatingDescent2024 extend the technique in and develop algorithms that can reach $\varepsilon$-stationary points within $\mathcal{O}(\varepsilon^{-3})$ iterations.
+Very recently, for problem (1) with additional assumptions that $h$ is smooth (but not necessarily convex) and $f(x,y)$ is linear with respect to $y$ for fixed $x$, the work xuRiemannianAlternatingDescent2024 extends the technique in heApproximationProximalGradient2024 and develops algorithms that can reach $\varepsilon$-stationary points within $\mathcal{O}(\varepsilon^{-3})$ iterations.
 It is worth noting that none of the methods in the aforementioned literature huangGradientDescentAscent2023; xuRiemannianAlternatingDescent2024; xuEfficientAlternatingRiemannian2024 can directly tackle problem (1) due to the following two reasons: (i) the convex function $h$ may be nonsmooth, and (ii) the function $f(x,y)$ is neither linear nor strongly concave with respect to $y$ for fixed $x$.
@@
 The main contributions of this paper are summarized below.
-First, we introduce the notions of optimization stationarity point as well as game stationarity for problem (1) and carefully study their relationships.
-Second, we propose a manifold proximal gradient descent ascent method for solving problem (1).
-At each iteration, this method performs one or multiple manifold proximal gradient descent steps to update $x$ and one proximal ascent step to update $y$.
-Third, we prove that the proposed method can find an $\varepsilon$-game-stationary point and an $\varepsilon$-optimization-stationary point within $O(\varepsilon^{-3})$ iterations.
-This iteration complexity result matches the state-of-the-art complexity analysis for single-loop algorithms in solving general nonsmooth NC-C optimization problems, where the constraint on $x$ is a compact convex set rather than a Riemannian manifold.
-Finally, we present numerical results on FSPCA and SSC to demonstrate the effectiveness of our proposed method.
+First, we introduce the notions of optimization stationarity as well as game stationarity for problem (1) and carefully study their relationships.
+Second, we propose two manifold proximal gradient descent ascent (MPGDA) algorithms for solving problem (1).
+The first algorithm, denoted by MPGDA-PA, performs one or multiple manifold proximal gradient descent steps to update $x$ and one proximal ascent step to update $y$ at each iteration.
+Note that the proximal ascent subproblem with respect to $y$ can generally be solved by FISTA (Beck2017First, Theorem 10.41) with a linear convergence rate.
+In particular, for the case of $f$ being linear with respect to $y$, addressing this subproblem reduces to solving a proximal problem in the form of (2b).
+The second algorithm, denoted by MPGDA-PGA, also performs one manifold proximal gradient descent step to update $x$ but uses a single proximal gradient ascent step to renew $y$ at each iteration.
+It is worth noting that the $y$-update in MPGDA-PGA requires only a single evaluation of the gradient $\nabla_{y}f$ and one step of solving the proximal problem (2b).
+This has much lower computational cost than that of MPGDA-PA in the case where $f$ is not linear with respect to $y$.
+Third, we prove that the MPGDA-PA algorithm can find an $\varepsilon$-game-stationary point of problem (1) within $\mathcal{O}(\varepsilon^{-3})$ outer iterations, while the MPGDA-PGA algorithm achieves this within $\mathcal{O}(\varepsilon^{-4})$ outer iterations.
+Finally, we illustrate the proposed algorithms via both analytic and simulated numerical examples.
 The rest of the paper is organized as follows.
@@
 Section 3 discusses the optimality condition of problem (1).
-Section 4 introduces the proposed MPGDA method and establish its iteration complexity.
+Section 4 introduces the proposed MPGDA-PA method and establishes its iteration complexity.
+Section 5 introduces the proposed MPGDA-PGA method and establishes its iteration complexity.
 Section 6 reports the numerical experiments.
 2 Preliminaries
-[…]
+In this section, we recall some basic notations and preliminary results which will be used in this paper.
+Let $\mathbb{E}$ represent a finite-dimensional Euclidean space with inner product $\langle\cdot,\cdot\rangle$ and the induced norm $\left\|\cdot\right\|$.
+For a linear operator $\mathcal{L}:\mathbb{E}_{1}\to\mathbb{E}_{2}$, let $\mathrm{rank}(\mathcal{L})$ denote the dimension of its image space $\{y\in\mathbb{E}_{2}\mid y=\mathcal{L}(x),~x\in\mathbb{E}_{1}\}$ and $\mathrm{ker}(\mathcal{L})$ denote its null space $\{x\in\mathbb{E}_{1}\mid\mathcal{L}(x)=0\}$ .
+Given a point $x\in\mathbb{E}$ and a set $C\subseteq\mathbb{E}$, let $\mathrm{dist}(x,C):=\inf_{y\in C}\|y-x\|$ and $\delta_{C}$ denote the indicator function associated with $C$.
+For a function $\varphi(x,y):\mathbb{E}_{1}\times\mathbb{E}_{2}\to\mathbb{R}$, let $\nabla_{x}\varphi(x,y)$ and $\nabla_{y}\varphi(x,y)$ respectively denote the partial gradients of $\varphi(x,y)$ with respect to $x$ and $y$ at the point $(x,y)$.
+Let $\mathcal{M}$ represent a Riemannian submanifold $\mathbb{E}_{1}$.
+For a function $\Phi:\mathcal{M}\to\mathbb{R}$, $\mathrm{grad}\Phi(x)$ denotes the Riemannian gradient of $\Phi$.
+For a function $\varphi:\mathcal{M}\times\mathbb{E}_{2}\to\mathbb{R}$, $\mathrm{grad}\varphi(\cdot,y)$ denotes the Riemannian gradient of $\varphi(\cdot,y)$ with $y$ fixed.
 2.1 Riemannian submanifold
@@
 Definition 1.
-(boumal2023intromanifolds, , Definition 3.10) Let $\mathbb{E}$ be a Euclidean space of dimension $d$.
-A nonempty subset $\mathcal{M}$ of $\mathbb{E}$ is called an embedded submanifold of $\mathbb{E}$ of dimension $n$ if either $n=d$ and $\mathcal{M}$ is open in $\mathbb{E}$ or $n=d-k$ for some $k\geq 1$ and, for each $x\in\mathcal{M}$, there exists a neighborhood $U$ of $x$ in $\mathbb{E}$ and a smooth function $\psi:U\to\mathbb{R}^{k}$ such that:
+(boumal2023intromanifolds, Definition 3.10) Let $\mathbb{E}$ be a Euclidean space of dimension $d$.
+A nonempty subset $\mathcal{M}$ of $\mathbb{E}$ is called an embedded submanifold of $\mathbb{E}$ of dimension $n$ if either $n=d$ and $\mathcal{M}$ is open in $\mathbb{E}$ or n=dkn=d-k for some $k\geq 1$ and, for each $x\in\mathcal{M}$, there exists a neighborhood $U$ of $x$ in $\mathbb{E}$ and a smooth function $\psi:U\to\mathbb{R}^{k}$ such that:
   1. 1.
@@
 Let $\mathcal{M}$ be an embedded submanifold of $\mathbb{E}$.
-The tangent space $T_{x}\mathcal{M}$ to $\mathcal{M}$ at $x\in\mathcal{M}$ is identified with $\mathbb{E}$ when $\mathcal{M}$ is open in $\mathbb{E}$; otherwise, $T_{x}\mathcal{M}=\mathrm{ker}(\mathrm{D}\psi(x))$ , where $\psi$ is a local defining function for $\mathcal{M}$ at $x$ (boumal2023intromanifolds, , Theorem 3.15).
+The tangent space $T_{x}\mathcal{M}$ to $\mathcal{M}$ at $x\in\mathcal{M}$ is identified with $\mathbb{E}$ when $\mathcal{M}$ is open in $\mathbb{E}$; otherwise, $T_{x}\mathcal{M}=\mathrm{ker}(\mathrm{D}\psi(x))$ , where $\psi$ is a local defining function for $\mathcal{M}$ at $x$ (boumal2023intromanifolds, Theorem 3.15).
 The embedded submanifold $\mathcal{M}$ is called a Riemannian submanifold of $\mathbb{E}$, when equipped with the Riemannian metric induced by the ambient space $\mathbb{E}$.
 Namely, for each $x\in\mathcal{M}$, the inner product on $T_{x}\mathcal{M}$ is defined as $\langle u,v\rangle_{x}:=\langle u,v\rangle$ for $u,~v\in T_{x}\mathcal{M}$.
-The normal space $N_{x}\mathcal{M}$ of a Riemannian submanifold $\mathcal{M}$ of $\mathbb{E}$ is the orthogonal complement of the tangent space $T_{x}\mathcal{M}$ in $\mathbb{E}$ (boumal2023intromanifolds, , Definition 5.48).
+The Riemannian gradient of a smooth function $\varphi:\mathbb{E}\to\mathbb{R}$ at a point $x\in\mathcal{M}$ is given by $\mathrm{grad}\varphi(x):=\mathrm{Proj}_{T_{x}\mathcal{M}}(\nabla\varphi(x))$, where $\mathrm{Proj}_{T_{x}\mathcal{M}}(\cdot)$ is the Euclidean projection operator onto $T_{x}\mathcal{M}$.
+The normal space $N_{x}\mathcal{M}$ of a Riemannian submanifold $\mathcal{M}$ of $\mathbb{E}$ is the orthogonal complement of the tangent space $T_{x}\mathcal{M}$ in $\mathbb{E}$ (boumal2023intromanifolds, Definition 5.48).
 The tangent bundle of $\mathcal{M}$ is the disjoint union of its tangent spaces : $T\mathcal{M}:=\{(x,v)\mid x\in\mathcal{M},~v\in T_{x}\mathcal{M}\}$.
-A retraction on a manifold $\mathcal{M}$ is a smooth mapping $R$ from the tangent bundle $T\mathcal{M}$ onto $\mathcal{M}$ such that, for any $(x,v)\in T_{x}\mathcal{M}$, its restriction $R_{x}:T_{x}\mathcal{M}\to\mathcal{M}$ satisfies (i) $R_{x}(0)=x$ and (ii) $\mathrm{D}R_{x}(0)[v]=v$ (boumal2023intromanifolds, , Definition 3.47).
-The concept of a retraction offers a practical link between manifold and its tangent bundle.
-The following properties of retractions are needed for convergence analysis in this paper.
+A retraction on a manifold $\mathcal{M}$ is a smooth mapping $R$ from the tangent bundle $T\mathcal{M}$ onto $\mathcal{M}$ such that, for any $(x,v)\in T_{x}\mathcal{M}$, its restriction $R_{x}:T_{x}\mathcal{M}\to\mathcal{M}$ satisfies (i) $R_{x}(0)=x$ and (ii) $\mathrm{D}R_{x}(0)[v]=v$ (boumal2023intromanifolds, Definition 3.47).
+The concept of a retraction offers a practical link between manifold and its tangent bundle. The following properties of retractions are needed for convergence analysis in this paper.
 Proposition 1.
-[…]
+(boumalGlobalRatesConvergence2019, Appendix B) Let $\mathcal{M}$ be a compact Riemannian submanifold of Euclidean space $\mathbb{E}$ equipped with norm $\left\|\cdot\right\|$.
+$R:T\mathcal{M}\to\mathcal{M}$ is a retraction.
+Then there exist constants $M_{1}>0$ and $M_{2}>0$ such that the following two inequalities hold for any $x\in\mathcal{M}$ and $v\in T_{x}{\mathcal{M}}$:
 $\left\|R_{x}(v)-x\right\|$ $\leq M_{1}\left\|v\right\|,$
@@
 2.2 Generalized subdifferential
-For a fucntion $\varphi:\mathbb{E}\to\mathbb{R}$, the usual one-sided directional derivative of $\varphi$ at $x\in\mathbb{E}$ along the direction $v\in\mathbb{E}$ is defined by
+For a function $\varphi:\mathbb{E}\to\mathbb{R}$, the usual one-sided directional derivative of $\varphi$ at $x\in\mathbb{E}$ along the direction $v\in\mathbb{E}$ is defined by
 $\varphi^{\prime}(x;v):=\lim_{t\downarrow 0}\frac{\varphi(x+tv)-\varphi(x)}{t}.$
-In contrast to the usual one-sided directional derivative, the (Clarke) generalized directional derivative dose not presuppose the existence of any limit (1983Optimization, , Section 2.1).
+In contrast to the usual one-sided directional derivative, the (Clarke) generalized directional derivative does not presuppose the existence of any limit (1983Optimization, Section 2.1).
 Specifically, the (Clarke) generalized directional derivative of $\varphi$ at $x\in\mathbb{E}$ along the direction $v\in\mathbb{E}$ is defined as
@@
 $\partial\varphi(x):=\{z\in\mathbb{E}\mid\langle z,v\rangle\leq\varphi^{\circ}(x;v),~\forall v\in\mathbb{E}\}.$
-The function $\varphi$ is said to be regular at $x$, provided that for all $v\in\mathbb{E}$, the usual one-sided directional derivative $\varphi^{\prime}(x;v)$ exists and $\varphi^{\prime}(x;v)=\varphi^{\circ}(x;v)$.
-If $\varphi_{1}$ and $\varphi_{2}$ are regular at $x$, then $\partial(\varphi_{1}+\varphi_{2})(x)=\partial\varphi_{1}(x)+\partial\varphi_{2}(x)$ (1983Optimization, , Definition 2.3.4).
-If $\varphi$ is a continuously differentiable function or a real-valued convex function, then it is regualr at any $x\in\mathbb{E}$ (1983Optimization, , Proposition 2.3.6).
-In addition, the indication function $\partial\delta_{\mathcal{M}}(x)$ of a Remannian submanifold $\mathcal{M}$ in $\mathbb{E}$ is regular at any $x\in\mathcal{M}$ (1983Optimization, , Theorem 2.4.7) and $\partial\delta_{\mathcal{M}}(x)=N_{x}\mathcal{M}$ (1983Optimization, , Proposition 2.4.12).
+If $\varphi:\mathbb{E}\to\mathbb{R}$ is convex, then for any $x\in\mathbb{E}$, $\partial f(x)$ coincides with the subdifferential at $x$ in the sense of convex analysis.
+A function $\varphi:\mathbb{E}\to\mathbb{R}$ is said to be regular at $x\in\mathbb{E}$, provided that for all $v\in\mathbb{E}$, the usual one-sided directional derivative $\varphi^{\prime}(x;v)$ exists and $\varphi^{\prime}(x;v)=\varphi^{\circ}(x;v)$ (1983Optimization, Definition 2.3.4).
+If $\varphi_{1}:\mathbb{E}\to\mathbb{R}$ and $\varphi_{2}:\mathbb{E}\to\mathbb{R}$ are locally Lipschitz continuous and regular at $x\in\mathbb{E}$, then $\varphi_{1}+\varphi_{2}$ are regular at $x$ and $\partial(\varphi_{1}+\varphi_{2})(x)=\partial\varphi_{1}(x)+\partial\varphi_{2}(x)$ (1983Optimization, Proposition 2.3.3, Corollary 3, Proposition 2.3.6).
+If $\varphi:\mathbb{E}\to\mathbb{R}$ is a continuously differentiable function or a real-valued convex function, then it is regualr at any $x\in\mathbb{E}$ (1983Optimization, Proposition 2.3.6).
+A function $\varphi:\mathbb{E}\to\mathbb{R}$ is said to be $\ell$-weakly convex if $\varphi(\cdot)+(\ell/2)\|\cdot\|^{2}$ is convex.
+It follows that the real-valued weakly convex function $\varphi$ is regular at any $x\in\mathbb{E}$.
+In addition, the indication function $\delta_{\mathcal{M}}(x)$ of a Remannian submanifold $\mathcal{M}$ in $\mathbb{E}$ is regular at any $x\in\mathcal{M}$ (1983Optimization, Theorem 2.4.7) and $\partial\delta_{\mathcal{M}}(x)=N_{x}\mathcal{M}$ (1983Optimization, Proposition 2.4.12).
 3 Optimality Conditions
 In this section, we introduce two important concepts of stationarity for problem (1), namely, the optimization stationarity and game stationarity, and examine their relationships.
-To this end, we first definite the value function $\Phi:\mathbb{E}\to\mathbb{R}$ by
+To this end, we first define the value function $\Phi:\mathbb{E}_{1}\to\mathbb{R}$ by
 $\Phi(x):=\max_{y\in S}f(x,y)-g(y).$ (9)
@@
 The next proposition concerns the value function $\Phi$.
-Recalled that, a function $\varphi:\mathbb{E}\to\mathbb{R}$ is said to be $\ell$-weakly convex if $\varphi(\cdot)+(\ell/2)\|\cdot\|^{2}$ is convex.
 Proposition 2.
-Let $y_{x}\in\arg\max_{y\in S}\{f(x,y)-g(y)\}$.
-The value function $\Phi$ is $L_{x}$-weakly convex and for all $x\in\mathcal{M}$, it holds that $\nabla_{x}f(x,y_{x})\in\partial\Phi(x)$.
-Moreover, if $f(x,y)-g(y)$ is strongly concave in $y$, then $\Phi$ is differentiable and $\nabla\Phi(x)=\nabla_{x}f(x,y_{x})$.
+The value function $\Phi$ defined in (9) is $L_{x}$-weakly convex and for all $x\in\mathbb{E}_{1}$, it holds that $\nabla_{x}f(x,y_{x})\in\partial\Phi(x)$ with $y_{x}\in\mathop{\mathrm{argmax}}_{y\in S}\{f(x,y)-g(y)\}$.
+Moreover, if f(x,y)g(y)f(x,y)-g(y) is strongly concave in $y$, then $\Phi$ is differentiable and $\nabla\Phi(x)=\nabla_{x}f(x,y_{x})$.
 Proof.
-Because $f(x,y)$ is $L_{x}$-smooth respect to $x$, $f(x,y)+\frac{L_{x}}{2}\left\|x\right\|^{2}-g(y)$ is convex in $x$ for any given $y\in S$.
+Because $f(x,y)$ is $L_{x}$-smooth with respect to $x$, $f(x,y)+\frac{L_{x}}{2}\left\|x\right\|^{2}-g(y)$ is convex in $x$ for any given $y\in S$.
 Thus, $\Phi(x)+\frac{L_{x}}{2}\left\|x\right\|^{2}=\max_{y\in S}f(x,y)+\frac{L_{x}}{2}\left\|x\right\|^{2}-g(y)$ is convex.
 Therefore, $\Phi(x)$ is $L_{x}$-weakly convex.
-Since $S$ is compact, according to Danskin’s theorem (bernhardTheoremDanskinApplication1995, , Theorem D1), $\nabla_{x}f(x,y_{x})\in\partial\Phi(x)$, and in particular, $\nabla\Phi(x)=\nabla_{x}f(x,y_{x})$ when $f(x,y)-g(y)$ is strongly concave in $y$.
+Since $S$ is compact, it follows from Danskin’s theorem (bernhardTheoremDanskinApplication1995, Theorem D1) that $\nabla_{x}f(x,y_{x})\in\partial\Phi(x)$, and in particular, $\nabla\Phi(x)=\nabla_{x}f(x,y_{x})$ when f(x,y)g(y)f(x,y)-g(y) is strongly concave in $y$.
 ∎
-In view of Proposition 2 and (1983Optimization, , Proposition 2.4.3), we see that any local minimizer $x^{*}$ of problem (9) must satisfy
+In view of Proposition 2 and (1983Optimization, Proposition 2.3.3, Proposition 2.3.6,Proposition 2.4.3), we see that any local minimizer $x^{*}$ of problem (10) must satisfy
 $0\in\partial\Phi(x^{*})+\partial h(x^{*})+\partial\delta_{\mathcal{M}}(x^{*}).$ (11)
-This naturally leads to the following definition of optimization stationarity.
+This naturally leads to the following definition of optimization stationarity, where the $\varepsilon$-optimization-stationarity is partly motivated by chenProximalGradientMethod2020 and perez2019subdifferential.
 Definition 2(Optimization stationarity).
@@
   2. (ii)
-    Let $\varepsilon>0$ be given. Inspired by perez2019subdifferential , we say that $x^{*}$ is an $\varepsilon$-optimization-stationary point of problem (1) if
-    $\mathrm{dist}(0,\partial_{\varepsilon}(\Phi+h)(x^{*})+\partial\delta_{\mathcal{M}}(x^{*}))<\varepsilon,$
-    with
-    $\partial_{\varepsilon}(\Phi+h)(x^{*}):=\{z\in\mathbb{E}\mid\langle z,v\rangle-\varepsilon\leq(\Phi+h)^{\circ}(x^{*};v),~{}\forall v\in\mathbb{E}\}.$
+    Let $\varepsilon>0$ be given. We say that $x^{*}$ is an $\varepsilon$-optimization-stationary point of problem (1) if there exists $u\in\mathbb{E}$ such that
+    $\max\{\|u\|,\mathrm{dist}(0,\partial_{\varepsilon}\Phi(x^{*}+u)+\partial h(x^{*}+u)+\partial\delta_{\mathcal{M}}(x^{*}))\}<\varepsilon,$
+    where
+    $\partial_{\varepsilon}\Phi(x):=\{z\in\mathbb{E}\mid\Phi(x)+\langle z,y-x\rangle-\varepsilon\leq\Phi(y)+\frac{L_{x}}{2}\|y-x\|^{2},~\forall y\in\mathbb{E}\}.$
 Now we consider defining the game stationarity of problem (1). It is routine to prove that any (local) saddle point $(x^{*},y^{*})$ of problem (1) satisfies
@@
 Suppose that $x^{*}$ is an optimization-stationary point of problem (1).
-By Proposition (2), $\Phi(x)+h(x)+L_{x}\|x-x^{*}\|^{2}$ is strongly convex.
-As $0\in\partial\delta_{\mathcal{M}}(x^{*})+\partial\Phi(x^{*})+\partial h(x^{*})$ and $\partial\delta_{\mathcal{M}}(x^{*})=N_{x^{*}}{\mathcal{M}}=\partial\delta_{T_{x^{*}}{\mathcal{M}}}(0)$, then $x^{*}$ is the unique minimum point of the problem
+By Proposition 2, $\Phi(x)+h(x)+L_{x}\|x-x^{*}\|^{2}$ is strongly convex.
+As $0\in\partial\delta_{\mathcal{M}}(x^{*})+\partial\Phi(x^{*})+\partial h(x^{*})$, and $\partial\delta_{\mathcal{M}}(x^{*})=N_{x^{*}}{\mathcal{M}}=\partial\delta_{T_{x^{*}}{\mathcal{M}}}(0)$, then $x^{*}$ is the unique minimum point of the problem
 $\min_{x\in A_{x^{*}}\mathcal{M}}\Phi(x)+h(x)+L_{x}\|x-x^{*}\|^{2}$
 with $A_{x^{*}}\mathcal{M}:=\{x\mid x=x^{*}+v,~v\in T_{x^{*}}{\mathcal{M}}\}$. Therefore, by Sion minimax theorem sionGeneralMinimaxTheorems1958, there exists $y^{*}\in S$ such that
-$f(x^{*},y^{*})+h(x^{*})-g(y^{*})$ $=\inf_{\begin{subarray}{c}x\in A_{x^{*}}{\mathcal{M}},\\ \|x-x^{*}\|\leq 1\end{subarray}}\sup_{y\in S}f(x,y)+h(x)-g(y)+L_{x}\|x-x^{*}\|^{2}$
+$f(x^{*},y^{*})+h(x^{*})-g(y^{*})$
+$=$ $\inf_{\begin{subarray}{c}x\in A_{x^{*}}{\mathcal{M}},\\ \|x-x^{*}\|\leq 1\end{subarray}}\sup_{y\in S}f(x,y)+h(x)-g(y)+L_{x}\|x-x^{*}\|^{2}$
 $=$ $\sup_{y\in S}\inf_{\begin{subarray}{c}x\in A_{x^{*}}{\mathcal{M}},\\ \|x-x^{*}\|\leq 1\end{subarray}}f(x,y)+h(x)-g(y)+L_{x}\|x-x^{*}\|^{2},$
@@
 $0\in\nabla_{y}f(x^{*},y^{*})-\partial g(y^{*})-\partial\delta_{S}(y^{*}).$
-Conversely, suppose that $(x^{*},y^{*})$ is a game stationary point of problem (1).
-Then, according to Proposition 2, $y^{*}=\arg\max_{y\in S}f(x^{*},y)-g(y)$ and $\nabla_{x}f(x^{*},y^{*})\in\partial\Phi(x^{*})$.
+Conversely, suppose that $(x^{*},y^{*})$ is a game-stationary point of problem (1).
+Then, according to Proposition 2, $y^{*}=\mathop{\mathrm{argmax}}_{y\in S}f(x^{*},y)-g(y)$ and $\nabla_{x}f(x^{*},y^{*})\in\partial\Phi(x^{*})$.
 Thus, $0\in\partial\Phi(x^{*})+\partial h(x^{*})+\partial\delta_{\mathcal{M}}(x^{*})$.
@@
 $\mathcal{G}^{\beta}(x,y):=\max\{\|\beta u\|,~\mathrm{dist}(0,\nabla_{y}f(x,y)-\partial g(y)-\partial\delta_{S}(y))\},$ (13)
-where $\beta>0$ and $u=\arg\min_{v\in T_{x}\mathcal{M}}\langle\nabla_{x}f(x,y),v\rangle+h(x+v)+\frac{\beta}{2}\|v\|^{2}$.
-This definition is motivated in part by the stationarity measure used in Euclidean minimax problem luHybridBlockSuccessive2020 .
+where $\beta>0$ and $u=\mathop{\mathrm{argmin}}_{v\in T_{x}\mathcal{M}}\langle\nabla_{x}f(x,y),v\rangle+h(x+v)+\frac{\beta}{2}\|v\|^{2}$.
+This definition is partly motivated by the stationarity measure used in the Euclidean minimax problem luHybridBlockSuccessive2020.
 It is straightforward to verify that $(x^{*},y^{*})$ is a game-stationary point of problem (1) if and only if there exists $\beta>0$ such that $\mathcal{G}^{\beta}(x^{*},y^{*})=0$.
@@
 Suppose that $(x^{*},y^{*})$ is an $\varepsilon$-game-stationary point of problem (1) with $\mathcal{G}^{\beta}(x^{*},y^{*})\leq\varepsilon$.
-Then $x^{*}$ is a $(C\varepsilon)$-optimization-stationary poin, where $C=(2\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}/\beta$ and $\sigma_{y}:=\sup_{y\in S}\left\|y\right\|$.
+Then, $x^{*}$ is a $(C\varepsilon)$-optimization-stationary point with
+$C=\max\{4\sigma_{y}^{2}+1+2L_{y}\sigma_{y}/\beta,1/\beta,1+L_{x}/\beta\},$
+where $L_{x},~L_{y},~\sigma_{y}$ are defined in Assumption 1.
 Proof.
-Suppose the point $(x^{*},y^{*})$ is an $\varepsilon$-game-stationary point of problem (1), then there exists $z$ such that $\|z\|=1$ and $\varepsilon z\in\nabla_{y}f(x^{*},y^{*})-\partial g(y^{*})-\partial\delta_{S}(y^{*})$.
-For $u=\arg\min_{v\in T_{x}\mathcal{M}}\langle\nabla_{x}f(x,y),v\rangle+h(x+v)+\frac{\beta}{2}\|v\|^{2}$, let
-$\Phi_{\varepsilon}^{u}(x):=$ $\max_{y\in S}f(x+u,y)-g(y)-\frac{\varepsilon}{2}\|y-(y^{*}-z)\|^{2}$
-$+\langle\nabla_{y}f(x^{*},y^{*})-\nabla_{y}f(x^{*}+u,y^{*}),y\rangle.$
-By Proposition 2 and the choice of $z$, we have $\nabla\Phi_{\varepsilon}^{u}(x^{*})=\nabla_{x}f(x^{*}+u,y^{*})$ and
-$\Phi_{\varepsilon}^{u}(x^{*})+\langle\nabla_{x}f(x^{*}+u,y),x-(x^{*}+u)\rangle\leq\Phi_{\varepsilon}^{u}(x-u)+\frac{L_{x}}{2}\|x-(x^{*}+u)\|^{2}.$
-Because
-$\Phi(x^{*}+u)-\varepsilon(2\sigma_{y}^{2}+1)-L_{y}\sigma_{y}\|u\|\leq\Phi_{\varepsilon}^{u}(x^{*})$
-and
-$\Phi_{\varepsilon}^{u}(x-u)\leq\Phi(x)+L_{y}\sigma_{y}\|u\|,$
-we arrive at
-$\Phi(x^{*}+u)+\langle\nabla_{x}f(x^{*}+u,y),x-(x^{*}+u)\rangle-\varepsilon(2\sigma_{y}^{2}+1)-2L_{y}\sigma_{y}\|u\|$
-$\leq$ $\Phi(x)+\frac{L_{x}}{2}\|x-(x^{*}+u)\|^{2}.$
-Since $h$ is convex and $L_{h}$-Lipschitz continuous, then
-$h(x^{*}+u)+\langle\xi,x-(x^{*}+u)\rangle\leq h(x)$
-for all $\xi\in\partial h(x^{*}+u)$. Therefore, by combing the above two inequalities, we obtain
-$\nabla_{x}f(x^{*}+u,y^{*})+\partial h(x^{*}+u)\subset\partial_{\hat{\varepsilon}}(\Phi+h)(x^{*}+u),$
-with $\hat{\varepsilon}=\varepsilon(2\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}\|u\|$.
-Furthermore, as $\|\mathcal{G}_{x}^{\beta}(x^{*},y^{*})\|\leq\varepsilon$, we can conclude that $x^{*}$ is a $(C\varepsilon)$-optimization-stationary point with $C=(2\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}/\beta$.
+Suppose the point $(x^{*},y^{*})$ is an $\varepsilon$-game-stationary point of problem (1), and let $u^{*}:=\arg\min_{v\in T_{x^{*}}\mathcal{M}}\langle\nabla_{x}f(x^{*},y^{*}),v\rangle+h(x^{*}+v)+\frac{\beta}{2}\|v\|^{2}$.
+Then, $\|\beta u^{*}\|\leq\varepsilon$, and there exists $z^{*}\in\mathbb{E}_{2}$ such that $\|z^{*}\|\leq 1$ and
+$\varepsilon z^{*}\in\nabla_{y}f(x^{*},y^{*})-\partial g(y^{*})-\partial\delta_{S}(y^{*}).$ (14)
+We first prove that
+$\nabla_{x}f(x^{*}+u^{*},y^{*})\in\partial_{\hat{\varepsilon}}\Phi(x^{*}+u^{*})\text{ with }\hat{\varepsilon}=\varepsilon(4\sigma_{y}^{2}+1)+2L_{y}\sigma_{y}\|u^{*}\|.$ (15)
+Consider the auxiliary function
+$\Phi_{\varepsilon}^{u^{*}}(x):=$ $\max_{y\in S}f(x+u^{*},y)-g(y)-\frac{\varepsilon}{2}\|y-(y^{*}-z^{*})\|^{2}$
+$+\langle\nabla_{y}f(x^{*},y^{*})-\nabla_{y}f(x^{*}+u^{*},y^{*}),y\rangle.$
+By (14), the point $y^{*}$ satisfies the first-order optimality condition for the maximization problem defining $\Phi^{u}_{\varepsilon}(x^{*})$.
+Hence, applying Danskin’s theorem (bernhardTheoremDanskinApplication1995, Theorem D1), we obtain $\nabla\Phi_{\varepsilon}^{u^{*}}(x^{*})=\nabla_{x}f(x^{*}+u^{*},y^{*})$.
+Following an argument similar to the proof of Proposition 2, we deduce that $\Phi_{\varepsilon}^{u^{*}}$ is $L_{x}$-weakly convex. Consequently, for any $x\in\mathbb{E}$,
+$\Phi_{\varepsilon}^{u^{*}}(x^{*})+\langle\nabla_{x}f(x^{*}+u^{*},y^{*}),x-(x^{*}+u^{*})\rangle$ (16)
+$\leq$ $\Phi_{\varepsilon}^{u^{*}}(x-u^{*})+\frac{L_{x}}{2}\|x-(x^{*}+u^{*})\|^{2}.$
+Recall that $\sigma_{y}$ is the upper bound of $S$ and $\nabla_{y}f(x,y)$ is $L_{y}$-Lipschitz continuous on $\mathcal{M}\times S$. Using $\|z^{*}\|\leq 1$, we obtain the estimates
+$\frac{\varepsilon}{2}\|y-(y^{*}-z^{*})\|^{2}\leq\varepsilon(4\sigma_{y}^{2}+1),$
+$\langle\nabla_{y}f(x^{*},y^{*})-\nabla_{y}f(x^{*}+u^{*},y^{*}),y\rangle\leq L_{y}\sigma_{y}\|u^{*}\|.$
+From the definitions of $\Phi$ and $\Phi_{\varepsilon}^{u^{*}}$, these inequalities yield
+$\Phi(x^{*}+u^{*})\leq\Phi_{\varepsilon}^{u^{*}}(x^{*})+\varepsilon(4\sigma_{y}^{2}+1)+L_{y}\sigma_{y}\|u^{*}\|,$ (17)
+and for any $x\in\mathbb{E}$,
+$\Phi_{\varepsilon}^{u^{*}}(x-u^{*})\leq\Phi(x)+L_{y}\sigma_{y}\|u^{*}\|.$ (18)
+Combining (16), (17) and (18), we obtain that for any $x\in\mathbb{E}$,
+$\Phi(x^{*}+u^{*})+\langle\nabla_{x}f(x^{*}+u^{*},y^{*}),x-(x^{*}+u^{*})\rangle-\varepsilon(4\sigma_{y}^{2}+1)-2L_{y}\sigma_{y}\|u^{*}\|$
+$\leq$ $\Phi(x)+\frac{L_{x}}{2}\|x-(x^{*}+u^{*})\|^{2}.$
+By the definition of the $\varepsilon$-subdifferential $\partial_{\varepsilon}\Phi(\cdot)$, this proves (15).
+On the other hand, from the definition of $u^{*}$, we have
+$\beta u^{*}-\nabla_{x}f(x^{*},y^{*})\in\partial h(x^{*}+u^{*})+\partial\delta_{\mathcal{M}}(x^{*}).$
+Combining this with (15) and recalling that $\nabla_{x}f(\cdot,y)$ is $L_{x}$-Lipschitz continuous, we deduce that
+$\mathrm{dist}(0,\partial_{\hat{\varepsilon}}\Phi(x^{*}+u^{*})+\partial h(x^{*}+u^{*})+\partial\delta_{\mathcal{M}}(x^{*}))<(\beta+L_{x})\|u^{*}\|.$
+Since $\|\beta u^{*}\|\leq\varepsilon$ by the $\varepsilon$-game-stationarity condition, it follows that $\hat{\varepsilon}\leq\varepsilon(4\sigma_{y}^{2}+1+2L_{y}\sigma_{y}/\beta)$ and $x^{*}$ is a $(C\varepsilon)$-optimization-stationary point with $C=\max\{4\sigma_{y}^{2}+1+2L_{y}\sigma_{y}/\beta,1/\beta,1+L_{x}/\beta\}$.
 ∎
-4 The proposed MPGDA algorithm
-In this section, we propose our MPGDA method for solving problem (1) and establish its iteration complexity.
+4 The proposed MPGDA-PA algorithm
+In this section, we propose the MPGDA-PA method for solving problem (1) and establish its iteration complexity.
 Our algorithm development is partially motivated by heApproximationProximalGradient2024; luHybridBlockSuccessive2020; xuRiemannianAlternatingDescent2024.
@@
 $Q_{k}(x):=h(x)+\Phi_{k}(x)$ (19)
-where $\Phi_{k}:\mathbb{E}_{1}\to\mathbb{R}$ is defined at $x\in\mathbb{E}_{1}$ by
+where $\Phi_{k}:\mathbb{E}_{1}\to\mathbb{R}$ is defined by
 $\Phi_{k}(x):=\max_{y\in S}\{f(x,y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}\}.$ (20)
 Here, $y_{k}$ denotes the iterate with respect to $y$ obtained in the previous iteration, $\gamma_{k}>0$ is the regularization parameter, and $\rho_{k}>0$ is the proximal parameter.
+Intuitively, the regularization term $\frac{\gamma_{k}}{2}\|y\|^{2}$ ensures strong concavity of the optimization problem in (9), while the proximal term $\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}$ stabilizes the update of $y$.
 Moreover, we choose $\gamma_{k}=\gamma_{0}/k^{1/3}$ with some $\gamma_{0}>0$ and assume the non-increasing sequence $\{\rho_{k}\}$ to be summable, i.e.,
 $\{\rho_{k}\}\in\mathcal{S}:=\{\{\rho_{k}\}\mid\sum_{k=0}^{\infty}\rho_{k}<+\infty,~~\rho_{k}\geq\rho_{k+1}>0,k=0,1,\dots~\}.$
-We further introduced the mapping $\bar{y}_{k}:\mathbb{E}_{1}\to\mathbb{E}_{2}$ defined by
-$\bar{y}_{k}(x):=\arg\max_{y\in S}\{f(x,y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}\}.$ (15)
-Then it holds that $Q_{k}(x)=F(x,\bar{y}_{k}(x))-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x)\|^{2}-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x)-y_{k}\|^{2}$ and in view of Danskin’s theorem (danskinTheoryMaxMinApplications1966, , Theorem D1), $\Phi_{k}$ is smooth on $\mathbb{E}_{1}$ and $\nabla\Phi_{k}(x)=\nabla_{x}f(x,\bar{y}_{k}(x))$.
-Now, at the $k$-th iteration, the proposed MPGDA method computes the new iterate by approximately solving the subproblem:
+We further introduce the mapping $\overline{y}_{k}:\mathbb{E}_{1}\to\mathbb{E}_{2}$ defined by
+$\overline{y}_{k}(x):=\mathop{\mathrm{argmax}}_{y\in S}\{f(x,y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}\}.$ (21)
+Then it holds that $\Phi_{k}(x)=f(x,\overline{y}_{k}(x))-\frac{\gamma_{k}}{2}\|\overline{y}_{k}(x)\|^{2}-\frac{\rho_{k}}{2}\|\overline{y}_{k}(x)-y_{k}\|^{2}$ and in view of Danskin’s theorem (danskinTheoryMaxMinApplications1966, Theorem D1), $\Phi_{k}$ is smooth on $\mathbb{E}_{1}$ and $\nabla\Phi_{k}(x)=\nabla_{x}f(x,\overline{y}_{k}(x))$.
+Now, at the $k$-th iteration, the proposed MPGDA-PA method computes the new iterate by approximately solving the subproblem:
 $\min_{x\in\mathcal{M}}Q_{k}(x).$ (22)
-Inspired by the structure of $Q_{k}$ and the ManPG algorithm (chenProximalGradientMethod2020, , Algorithm 4.1), we perform Riemannian manifold proximal gradient descent steps for $T_{k}$ times to generate $x_{k+1}$ with $T_{k}$ being an integer in $[1,\overline{T}]$ where $\overline{T}$ is a predetermined positive integer.
-Then $y_{k+1}$ is updated as $\bar{y}_{k}(x_{k+1})$ with $\bar{y}_{k}$ defined in (15).
-Notice that the update of $y$ can also be interpreted as performing a proximal ascent step on the function $f(x_{k+1},y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}$.
-We formally presented the proposed method in Algorithm 1.
-Algorithm 1 Manifold Proximal Gradient Descent Ascent (MPGDA) for problem (1)
-0:  $c_{1}>0$, $\eta\in(0,1)$, $\gamma_{0}>0$, $x_{0}\in\mathcal{M}$, $y_{0}\in\mathrm{dom}(g)\cap S$, $0<l_{\min}<l_{\max}$, $\{\rho_{k}\}\in\mathcal{S}$, $\{\gamma_{k}=\gamma_{0}/k^{1/3}:k\in N\}$, $\{T_{k}\}\subset[1,\overline{T}]$;
-1:  for $k=0,\dots$ do
-2:     Set $x_{k,0}=x_{k}$,
+Inspired by the structure of $Q_{k}$ and the ManPG algorithm (chenProximalGradientMethod2020, Algorithm 4.1), we perform Riemannian manifold proximal gradient descent steps for $T_{k}$ times to generate $x_{k+1}$ with $T_{k}$ being an integer in $1,\overline{T}]$ where $\overline{T}$ is a predetermined positive integer.
+Then $y_{k+1}$ is updated as $\overline{y}_{k}(x_{k+1})$ with $\overline{y}_{k}$ defined in (21).
+Note that the update of $y$ can also be interpreted as performing a proximal ascent step for the function $f(x_{k+1},y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}$ on $S$.
+We formally present the proposed method in Algorithm 1.
+Algorithm 1 Manifold Proximal Gradient Descent Ascent (MPGDA-PA) for problem (1)
+[…]
+1:for $k=0,1,\dots$ do
+2:  Set $x_{k,0}=x_{k}$.
 3:  for $i=0,\dots,T_{k}-1$ do
@@
 5:   Compute
-$v_{k,i}:=\arg\min_{v\in T_{x_{k,i}}\mathcal{M}}\langle\nabla_{x}f(x_{k,i},\bar{y}_{k}(x_{k,i})),v\rangle+h(x_{k,i}+v)+\frac{\beta_{k,i}}{2}\left\|v\right\|^{2},$ (17)
+$v_{k,i}:=\mathop{\mathrm{argmin}}_{v\in T_{x_{k,i}}\mathcal{M}}\langle\nabla_{x}f(x_{k,i},\overline{y}_{k}(x_{k,i})),v\rangle+h(x_{k,i}+v)+\frac{\beta_{k,i}}{2}\left\|v\right\|^{2},$ (23)
 6:   Find the smallest non-negative integer $j$ such that
 $Q_{k}(R_{x_{k,i}}(\eta^{j}v_{k,i}))\leq Q_{k}(x_{k,i})-c_{1}\eta^{j}\beta_{k,i}\|v_{k,i}\|^{2}+2\rho_{k}\sigma_{y}^{2}$ (24)
-and update $x_{k,i+1}=R_{x_{k,i}}(\eta^{j}v_{k,i})$
+and update $x_{k,i+1}=R_{x_{k,i}}(\eta^{j}v_{k,i})$.
 7:  end for
-8:     $x_{k+1}=x_{k,T_{k}},~{}y_{k+1}=\bar{y}_{k}(x_{k+1})$.
+8:  $x_{k+1}=x_{k,T_{k}},~y_{k+1}=\overline{y}_{k}(x_{k+1})$.
 9:end for
 10:return $(x_{k+1},y_{k+1})$.
-Before proceeding, we make two remarks on the iteration procedure of MPGDA algorithm.
-First, given $x_{k,i}\in\mathcal{M}$, MPGDA algorithm computes a descent direction $v_{k,i}$ of $Q_{k}$ restricted to the tangent space $T_{x_{k,i}}\mathcal{M}$ via tackling the linear constrained strongly convex optimization problem (17), which can usually be efficiently solved, see (li2024proximal, , Section 4), for example.
-Second, in practice, we make some choices of parameters $l_{k,i}$ and $\rho_{k}$, which enhance the efficiency of MPGDA algorithm in our test and also satisfy $l_{k,i}\in[l_{\min},l_{\max}]$ and $\rho_{k}\in S$.
-For parameters $l_{k,i}$, we adopt the Riemannian Barzilai-Borwein (BB) stepsize iannazzo2018riemannian ; 2013AWen as follows
-$l_{k,i}=\begin{cases}\min\{\max\{l_{\min},(\rho_{k}+\gamma_{k})|\frac{\langle\Delta X,\Delta G\rangle}{\|\Delta X\|^{2}}|)\},l_{\max}\},&if~{}\Delta X\neq 0,\\ l_{\max},&if~{}\Delta X=0,\end{cases}$
-where $\Delta X=X_{k,i}-X_{k,i-1}$ and $\Delta G=\mathrm{grad}\Phi_{k}(x_{k,i})-\mathrm{grad}\Phi_{k}(x_{k,i-1})$.
+Before proceeding, we make some remarks on the iteration procedure of the MPGDA-PA algorithm.
+First, given $x_{k,i}\in\mathcal{M}$, the MPGDA-PA algorithm computes a descent direction $v_{k,i}$ of $Q_{k}$ restricted to the tangent space $T_{x_{k,i}}\mathcal{M}$ via tackling the linear constrained strongly convex optimization problem (23), which can usually be efficiently solved, see (li2024proximal, Section 4), for example.
+Second, for all $x\in\mathcal{M}$, $\overline{y}_{k}(x)$ defined in (21) can be solved by applying FISTA Beck2017First to problem (20), which converges linearly due to the strong convexity and smoothness of
+$\frac{\gamma_{k}}{2}\|\cdot\|^{2}+\frac{\rho_{k}}{2}\|\cdot-y_{k}\|^{2}-f(x,\cdot).$
+In particular, if $f(x,y)$ is linear with respect to $y$ for fixed $x$, i.e., $f(x,y)=\langle\mathcal{A}(x),y\rangle$, then $\overline{y}_{k}(x)$ can be further simplified into a proximal problem like (2b) as follows:
+$\overline{y}_{k}(x)=\mathop{\mathrm{argmax}}_{y\in S}\left\{-g(y)-\frac{\gamma_{k}+\rho_{k}}{2}\|y-\frac{\rho_{k}y_{k}+\mathcal{A}(x)}{\rho_{k}+\gamma_{k}}\|^{2}\right\}.$
+Third, in practice, we make some choices of parameters $l_{k,i}$ and $\rho_{k}$, which enhance the efficiency of the MPGDA-PA algorithm in our test and also satisfy $l_{k,i}\in[l_{\min},l_{\max}]$ and $\rho_{k}\in S$.
+Motivated by the Riemannian Barzilai-Borwein (BB) stepsize iannazzo2018riemannian; 2013AWen, we choose $l_{k,i}$ as follows:
+$l_{k,i}=\begin{cases}\min\{\max\{l_{\min},(\rho_{k}+\gamma_{k})|\frac{\langle\Delta X,\Delta R\rangle}{\|\Delta X\|^{2}}|)\},l_{\max}\},&if~\Delta X\neq 0,\\ l_{\max},&if~\Delta X=0,\end{cases}$
+where $\Delta X=X_{k,i}-X_{k,i-1}$ and $\Delta R=\mathrm{grad}\Phi_{k}(x_{k,i})-\mathrm{grad}\Phi_{k}(x_{k,i-1})$.
 Moreover, inspired by xuRiemannianAlternatingDescent2024, for given constant $\theta>1$, $\tau_{1}\in(0,1)$ and $\tau_{2}\in(0,1)$, we set $\rho_{0}=\xi_{0}$ and for $k\geq 1$,
@@
 $\delta_{k}=\|\gamma_{k-1}y_{k}+\rho_{k-1}(y_{k}-y_{k-1})\|_{\infty}.$
-In what follows, we conduct convergence analysis for MPGDA algorithm. First, we present a technical lemma that guarantees the boundedness of the mapping $\bar{y}_{k}$.
+In what follows, we conduct convergence analysis for the MPGDA-PA algorithm. First, we present a technical lemma that guarantees the boundedness of the mapping $\overline{y}_{k}$.
 Lemma 1.
-Let the mapping $\bar{y}_{k}:\mathbb{E}_{1}\to\mathbb{E}_{2}$ be defined in (15) for $k\in N$. Then for all $x$, $\bar{x}\in\mathbb{E}_{1}$ and $y\in\mathbb{E}_{2}$ , it holds that
+Let the mapping $\overline{y}_{k}:\mathbb{E}_{1}\to\mathbb{E}_{2}$ be defined in (21) for $k\in\mathbb{N}$. Then for all $x$, $\bar{x}\in\mathbb{E}_{1}$ and $y\in\mathbb{E}_{2}$ , it holds that
 $f(x,y)-g(y)$
-$\leq f(x,\bar{y}_{k}(\bar{x}))-g(\bar{y}_{k}(\bar{x}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(\bar{x})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(\bar{x})\|^{2}+\frac{\gamma_{k}}{2}\|y\|^{2}$
+$\leq f(x,\overline{y}_{k}(\bar{x}))-g(\overline{y}_{k}(\bar{x}))-\frac{\rho_{k}}{2}\|\overline{y}_{k}(\bar{x})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\overline{y}_{k}(\bar{x})\|^{2}+\frac{\gamma_{k}}{2}\|y\|^{2}$
 $+\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}+\frac{L_{y}^{2}}{2(\rho_{k}+\gamma_{k})}\|x-\bar{x}\|^{2}.$
 Proof.
-According to the definition of $\bar{y}_{k}(\bar{x})$, we have
-$0\in\nabla_{y}f(\bar{x},\bar{y}_{k}(\bar{x}))-\partial g(\bar{y}_{k}(\bar{x}))-\partial\delta_{S}(\bar{y}_{k}(\bar{x}))-\rho_{k}(\bar{y}_{k}(\bar{x})-y_{k})-\gamma_{k}\bar{y}_{k}(\bar{x}).$
-By subgradient inequality, we obtain
+By the definition of $\overline{y}_{k}(\bar{x})$, we have
+$0\in\nabla_{y}f(\bar{x},\overline{y}_{k}(\bar{x}))-\partial g(\overline{y}_{k}(\bar{x}))-\partial\delta_{S}(\overline{y}_{k}(\bar{x}))-\rho_{k}(\overline{y}_{k}(\bar{x})-y_{k})-\gamma_{k}\overline{y}_{k}(\bar{x}).$
+This and the concavity of $f(x,\cdot)-g(\cdot)$ yield that
 $f(x,y)-g(y)$
-$\leq f(x,\bar{y}_{k}(\bar{x}))-g(\bar{y}_{k}(\bar{x}))+\langle\nabla_{y}f(x,\bar{y}_{k}(\bar{x}))-\nabla_{y}f(\bar{x},\bar{y}_{k}(\bar{x})),y-\bar{y}_{k}(\bar{x})\rangle$
-$+\langle\rho_{k}(\bar{y}_{k}(\bar{x})-y_{k})+\gamma_{k}\bar{y}_{k}(\bar{x}),y-\bar{y}_{k}(\bar{x})\rangle.$
-Since $2\langle a,a-b\rangle=\left\|a\right\|^{2}+\left\|a-b\right\|^{2}-\left\|b\right\|^{2}$,by Cauchy-Schwarz inequality, we arrive at
+$\leq f(x,\overline{y}_{k}(\bar{x}))-g(\overline{y}_{k}(\bar{x}))+\langle\nabla_{y}f(x,\overline{y}_{k}(\bar{x}))-\nabla_{y}f(\bar{x},\overline{y}_{k}(\bar{x})),y-\overline{y}_{k}(\bar{x})\rangle$
+$+\langle\rho_{k}(\overline{y}_{k}(\bar{x})-y_{k})+\gamma_{k}\overline{y}_{k}(\bar{x}),y-\overline{y}_{k}(\bar{x})\rangle.$
+Since $2\langle a,a-b\rangle=\left\|a\right\|^{2}+\left\|a-b\right\|^{2}-\left\|b\right\|^{2}$ and $\nabla_{y}f(x,y)$ is $L_{y}$-Lipschitz continuous, it follows that,
 $f(x,y)-g(y)$
-$\leq f(x,\bar{y}_{k}(\bar{x}))-g(\bar{y}_{k}(\bar{x}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(\bar{x})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(\bar{x})\|^{2}+\frac{\gamma_{k}}{2}\|y\|^{2}$
+$\leq f(x,\overline{y}_{k}(\bar{x}))-g(\overline{y}_{k}(\bar{x}))-\frac{\rho_{k}}{2}\|\overline{y}_{k}(\bar{x})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\overline{y}_{k}(\bar{x})\|^{2}+\frac{\gamma_{k}}{2}\|y\|^{2}$
 $+\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}+\frac{L_{y}^{2}}{2(\rho_{k}+\gamma_{k})}\|x-\bar{x}\|^{2}.$
-Next, we prove that the descent condition (18) must be satisfied after a unified finite number of iterations for all $k\in N$ and $i\in[0,T_{k}-1]$, which ensures that the proposed Algorithm 1 is well-defined.
+Next, we prove that the descent condition (24) must be satisfied after a unified finite number of iterations for all $k\in\mathbb{N}$ and $i\in[0,T_{k}-1]$, which ensures that the proposed Algorithm 1 is well-defined.
 Proposition 5.
-Consider Algorithm 1. There exists a constant $\bar{\alpha}>0$ such that for all $k\in N$ and $0<\alpha\leq\min\{1,~{}\bar{\alpha}\}$, it holds that
+Consider Algorithm 1. There exists a constant $\bar{\alpha}>0$ such that for all $k\in\mathbb{N}$ and $0<\alpha\leq\min\{1,~\bar{\alpha}\}$, it holds that
 $Q_{k}(R_{x_{k,i}}(\alpha v_{k,i}))\leq Q_{k}(x_{k,i})-\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}+2\rho_{k}\sigma_{y}^{2}$
@@
 Proof.
-By (chenProximalGradientMethod2020, , Lemma 5.2), for any $0<\alpha\leq\min\{1,~{}\frac{\beta_{k,i}}{2M_{2}(L_{h}+G)+M_{1}^{2}L_{x}}\}$,
-$f(x_{\alpha},\bar{y}_{k}(x_{k,i}))+h(x_{\alpha})\leq f(x_{k,i},\bar{y}_{k}(x_{k,i}))+h(x_{k,i})-\frac{\alpha\beta_{k,i}}{2}\|v_{k,i}\|^{2},$
-where $x_{\alpha}=R_{x_{k,i}}(\alpha v_{k,i})$ and $G=\sup_{x\in\mathcal{M},y\in\mathrm{dom}(g)\cap S}\|\nabla_{x}F(x,y)\|$.
+Let $G:=\sup\{\|\nabla_{x}f(x,y)\|\mid x\in\mathcal{M},y\in S\}<\infty.$ Assume that $0<\alpha\leq\min\{1,~\frac{\beta_{k,i}}{2M_{2}(L_{h}+G)+M_{1}^{2}L_{x}}\}$ with $M_{1}>0,~M_{2}>0$ being defined in Proposition 1 and $x_{\alpha}=R_{x_{k,i}}(\alpha v_{k,i})$.
+Then, following an argument similar to the proof of (chenProximalGradientMethod2020, Lemma 5.2), we deduce that
+$f(x_{\alpha},\overline{y}_{k}(x_{k,i}))+h(x_{\alpha})\leq f(x_{k,i},\overline{y}_{k}(x_{k,i}))+h(x_{k,i})-\frac{\alpha\beta_{k,i}}{2}\|v_{k,i}\|^{2}.$
 According to Lemma 1, we have
-$f(x_{\alpha},\bar{y}_{k}(x_{\alpha}))-g(\bar{y}_{k}(x_{\alpha}))$
-$\leq f(x_{\alpha},\bar{y}_{k}(x_{k,i}))-g(\bar{y}_{k}(x_{k,i}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{k,i})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{k,i})\|^{2}+\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{\alpha})\|^{2}$
-$+\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{\alpha})-y_{k}\|^{2}+\frac{L_{y}^{2}}{2(\rho_{k}+\gamma_{k})}\|x_{\alpha}-x_{k,i}\|^{2}.$
-By summing these inequalities with the inequality
+$f(x_{\alpha},\overline{y}_{k}(x_{\alpha}))-g(\overline{y}_{k}(x_{\alpha}))$
+$\leq f(x_{\alpha},\overline{y}_{k}(x_{k,i}))-g(\overline{y}_{k}(x_{k,i}))-\frac{\rho_{k}}{2}\|\overline{y}_{k}(x_{k,i})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\overline{y}_{k}(x_{k,i})\|^{2}$
+$+\frac{\gamma_{k}}{2}\|\overline{y}_{k}(x_{\alpha})\|^{2}+\frac{\rho_{k}}{2}\|\overline{y}_{k}(x_{\alpha})-y_{k}\|^{2}+\frac{L_{y}^{2}}{2(\rho_{k}+\gamma_{k})}\|x_{\alpha}-x_{k,i}\|^{2}.$
+By summing the above two inequalities and the inequality
 $\|x_{\alpha}-x_{k,i}\|\leq\alpha M_{1}\|v_{k,i}\|$
-in Proposition 1, we derive the following inequality:
-$F(x_{\alpha},\bar{y}_{k}(x_{\alpha}))-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{\alpha})\|^{2}$
-$\leq F(x_{k,i},\bar{y}_{k}(x_{k,i}))-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{k,i})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x_{k,i})\|^{2}$
-$+\frac{\rho_{k}}{2}\|\bar{y}_{k}(x_{\alpha})-y_{k}\|^{2}-(\frac{\alpha\beta_{k,i}}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2(\rho_{k}+\gamma_{k})})\|v_{k,i}\|^{2}.$
-Since $\beta_{k,i}=l_{k,i}/(\rho_{k}+\gamma_{k})$, for any $0<\alpha\leq\frac{l_{k,i}}{2M_{1}^{2}L_{y}^{2}}$, we have
-$(\frac{\alpha\beta_{k,i}}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2(\rho_{k}+\gamma_{k})})\|v_{k,i}\|^{2}$ $=\beta_{k,i}(\frac{\alpha}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2l_{k,i}})\|v_{k,i}\|^{2}\geq\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}.$
-Note that $Q_{k}(x)=F(x,\bar{y}_{k}(x))-\frac{\gamma_{k}}{2}\|\bar{y}_{k}(x)\|^{2}-\frac{\rho_{k}}{2}\|\bar{y}_{k}(x)-y_{k}\|^{2}$, then for any $0<\alpha\leq\min\{1,\frac{l_{\min}}{(\rho_{0}+\gamma_{0})(2M_{2}(L_{h}+G)+M_{1}^{2}L_{x})},\frac{l_{\min}}{2M_{1}^{2}L_{y}^{2}}\}$, we have
-$Q_{k}(x_{\alpha})\leq Q_{k}(x_{k,i})-\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}+2\rho_{k}\sigma_{y}^{2}$
-for all $k\geq 0$. ∎
-We are now ready to establish the iteration complexity for MPGDA algorithm. Let $\varepsilon>0$ be a given target accuracy, we define
-$T(\varepsilon):=\min\{k\mid\mathcal{G}^{\beta_{k,0}}(x_{k,0},\bar{y}_{k}(x_{k,0}))<\varepsilon\}$
+from Proposition 1, we derive the following inequality:
+$F(x_{\alpha},\overline{y}_{k}(x_{\alpha}))-\frac{\gamma_{k}}{2}\|\overline{y}_{k}(x_{\alpha})\|^{2}$ (26)
+$\leq F(x_{k,i},\overline{y}_{k}(x_{k,i}))-\frac{\rho_{k}}{2}\|\overline{y}_{k}(x_{k,i})-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|\overline{y}_{k}(x_{k,i})\|^{2}$
+$+\frac{\rho_{k}}{2}\|\overline{y}_{k}(x_{\alpha})-y_{k}\|^{2}-(\frac{\alpha\beta_{k,i}}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2(\rho_{k}+\gamma_{k})})\|v_{k,i}\|^{2}.$
+Since $\beta_{k,i}=l_{k,i}/(\rho_{k}+\gamma_{k})$, for any $0<\alpha\leq\frac{l_{k,i}}{2M_{1}^{2}L_{y}^{2}}$, it holds that
+$(\frac{\alpha\beta_{k,i}}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2(\rho_{k}+\gamma_{k})})\|v_{k,i}\|^{2}$ $=\beta_{k,i}(\frac{\alpha}{2}-\frac{\alpha^{2}M_{1}^{2}L_{y}^{2}}{2l_{k,i}})\|v_{k,i}\|^{2}\geq\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}.$ (27)
+Combining (LABEL:ineq-F-gammak), (27), and the definition of $Q_{k}(x)$, we obtain that for all $k\geq 0$ and any $0<\alpha\leq\min\{1,\bar{\alpha}\}$,
+$Q_{k}(x_{\alpha})\leq Q_{k}(x_{k,i})-\frac{\alpha}{4}\beta_{k,i}\|v_{k,i}\|^{2}+2\rho_{k}\sigma_{y}^{2},$
+where $\bar{\alpha}=\min\{\frac{l_{\min}}{(\rho_{0}+\gamma_{0})(2M_{2}(L_{h}+G)+M_{1}^{2}L_{x})},\frac{l_{\min}}{2M_{1}^{2}L_{y}^{2}}\}$. ∎
+By Proposition 5, the backtracking line search procedure in Algorithm 1 will terminate after a finite number of steps. We present this result in the following proposition.
+Proposition 6.
+For all $k\geq 0$, the condition (24) in Algorithm 1 is satisfied within at most $J_{1}$ backtracking steps, where
+[…]
+We are now ready to establish the iteration complexity for Algorithm 1. Let $\varepsilon>0$ be a given target accuracy, we define
+$T(\varepsilon):=\min\{k\mid\mathcal{G}^{\beta_{k}}(x_{k},y_{k})<\varepsilon\}$ (28)
 where $\mathcal{G}^{\beta}(\cdot,\cdot)$ is defined in (13).
-In other words, $T(\varepsilon)$ stands for the minimal number of outer iterations needed for MPGDA algorithm to obtain a $\varepsilon$-game-stationary point.
-We further define two constants $\overline{Q}$ and $\underline{Q}$ as
-$\overline{Q}$ $:=\sup\{{F(x,y)\mid x\in\mathcal{M},y\in\mathrm{dom}(g)\cap S}\},$
-$\underline{Q}$ $:=\inf\{F(x,y)-\frac{\rho_{0}}{2}\|y-y_{0}\|^{2}-\frac{\gamma_{0}}{2}\|y\|^{2}\mid x\in\mathcal{M},y\in\mathrm{dom}(g)\cap S\}.$
+In other words, $T(\varepsilon)$ stands for the minimal number of outer iterations needed for Algorithm 1 to obtain a $\varepsilon$-game-stationary point.
+We further introduce two constants $\overline{Q}$ and $\underline{Q}$ as follows
+$\overline{Q}$ $:=\sup\{{F(x,y)\mid x\in\mathcal{M},y\in S}\},$
+$\underline{Q}$ $:=\inf\{F(x,y)-\frac{\rho_{0}}{2}\|y-y_{0}\|^{2}-\frac{\gamma_{0}}{2}\|y\|^{2}\mid x\in\mathcal{M},y\in S\}.$
 Theorem 4.1.
-Given $\varepsilon>0$, let $\{(x_{k},y_{k})\}$ be generated by Algorithm 1. Then, there exists a positive $\bar{\alpha}>0$ such that the following inequality holds:
-$T(\varepsilon)\leq\frac{\max\{(2\rho_{0}+\gamma_{0})\sigma_{y},C\}}{\varepsilon^{3}}$
+Given $\varepsilon>0$, let $\{(x_{k},y_{k})\}$ be the sequence generated by Algorithm 1. Then, the following inequality holds:
+$T(\varepsilon)\leq\frac{(\max\{(2\rho_{0}+\gamma_{0})\sigma_{y},C\})^{3}}{\varepsilon^{3}}+1$
 with
-$C=\frac{\sqrt{2(\overline{Q}-\underline{Q}+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2})l_{\max}}}{\sqrt{3\bar{\alpha}\eta\gamma_{0}}}.$
-Here, all the parameters $\rho_{0},~{}\gamma_{0},~{}\sigma_{y},~{}\overline{Q},~{}\underline{Q},~{}\overline{T},~{}S_{\rho},~{}l_{\max},~{}\eta$ are defined as in the relevant context of the paper.
+$C=\frac{\sqrt{2(\overline{Q}-\underline{Q}+\frac{\gamma_{0}}{2}\sigma_{y}^{2}+2(\overline{T}+1)S_{\rho}\sigma_{y}^{2})l_{\max}}}{\sqrt{3c_{1}\eta^{J_{1}}\gamma_{0}}}.$
+Here, the constants $c_{1},~\eta,~\rho_{0},~\gamma_{0},~\overline{T}$, and $l_{\max}$ are parameters of Algorithm 1, $J_{1}$ is defined in Proposition 6, and $S_{\rho}:=\sum_{k=0}^{\infty}\rho_{k}<\infty$.
 Proof.
-Let $\beta_{k}:=\beta_{k,0},~{}v_{k}:=v_{k,0}$. Since $\gamma_{k+1}\leq\gamma_{k}$ and $\rho_{k+1}\leq\rho_{k}$, by the definition of $Q_{k}$, we have
-$Q_{k+1}(x_{k+1})$ $=h(x_{k+1})+\max_{y\in S}f(x_{k+1},y)-g(y)-\frac{\gamma_{k+1}}{2}\|y\|^{2}-\frac{\rho_{k+1}}{2}\|y-y_{k+1}\|^{2}$
-$\leq h(x_{k,T_{k}})+\max_{y\in S}f(x_{k,T_{k}},y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}+4\rho_{k}\sigma_{y}^{2}$
-$=Q_{k}(x_{k,T_{k}})+4\rho_{k}\sigma_{y}^{2}.$
-Furthermore, according to Proposition 5, there exists $\bar{\alpha}>0$ such that
-$Q_{k+1}(x_{k+1})\leq Q_{k}(x_{k})+2(T_{k}+2)\rho_{k}\sigma_{y}^{2}-\bar{\alpha}\eta\beta_{k}\|v_{k}\|^{2}.$
+Let $\beta_{k}:=\beta_{k,0},~v_{k}:=v_{k,0}$. Since $\gamma_{k+1}\leq\gamma_{k}$, $\rho_{k+1}\leq\rho_{k}$, and $x_{k+1}=x_{k,T_{k}}$, by the definition of $Q_{k}$ and $\sigma_{y}$, we have
+$Q_{k+1}(x_{k+1})$
+$=$ $h(x_{k+1})+\max_{y\in S}f(x_{k+1},y)-g(y)-\frac{\gamma_{k+1}}{2}\|y\|^{2}-\frac{\rho_{k+1}}{2}\|y-y_{k+1}\|^{2}$
+$\leq$ $h(x_{k,T_{k}})+\max_{y\in S}f(x_{k,T_{k}},y)-g(y)-\frac{\gamma_{k}}{2}\|y\|^{2}-\frac{\rho_{k}}{2}\|y-y_{k}\|^{2}$
+$+\frac{\gamma_{k}-\gamma_{k+1}}{2}\sigma_{y}^{2}+2\rho_{k}\sigma_{y}^{2}$
+$=$ $Q_{k}(x_{k,T_{k}})+\frac{\gamma_{k}-\gamma_{k+1}}{2}\sigma_{y}^{2}+2\rho_{k}\sigma_{y}^{2}.$
+Combining this with (24) and Proposition 6, we obtain that
+$Q_{k+1}(x_{k+1})\leq Q_{k}(x_{k})+2(T_{k}+1)\rho_{k}\sigma_{y}^{2}+\frac{\gamma_{k}-\gamma_{k+1}}{2}\sigma_{y}^{2}-c_{1}\eta^{J_{1}}\beta_{k}\|v_{k}\|^{2}.$
 Summing the above inequality over $k=K_{1},K_{1}+1,\dots,K$ yields
-$\sum_{k=K_{1}}^{K}\bar{\alpha}\eta\beta_{k}\|v_{k}\|^{2}\leq{Q}_{K_{1}}(x_{K_{1}})-{Q}_{K}(x_{K})+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2}$ (20)
-with $S_{\rho}=\sum_{k=0}^{\infty}\rho_{k}<\infty$.
-Since $\beta_{k}\leq\frac{l_{\max}}{\gamma_{0}k^{-1/3}}$, then
-$\sum_{k=K_{1}}^{K}\bar{\alpha}\eta\beta_{k}\|v_{k}\|^{2}\geq(\frac{\bar{\alpha}\eta\gamma_{0}}{l_{\max}}\sum_{k=K_{1}}^{K}k^{1/3})\min_{K_{1}\leq k\leq K}\|\beta_{k}v_{k}\|^{2}.$ (21)
+$\sum_{k=K_{1}}^{K}c_{1}\eta^{J_{1}}\beta_{k}\|v_{k}\|^{2}\leq{Q}_{K_{1}}(x_{K_{1}})-{Q}_{K}(x_{K+1})+\frac{\gamma_{K_{1}}-\gamma_{K}}{2}\sigma_{y}^{2}+2(\overline{T}+1)S_{\rho}\sigma_{y}^{2}$ (29)
+with $S_{\rho}=\sum_{k=0}^{\infty}\rho_{k}<\infty$. Since $\beta_{k}\leq\frac{l_{\max}}{\gamma_{0}k^{-1/3}}$, then
+$\sum_{k=K_{1}}^{K}c_{1}\eta^{J_{1}}\beta_{k}\|v_{k}\|^{2}\geq(\frac{c_{1}\eta^{J_{1}}\gamma_{0}}{l_{\max}}\sum_{k=K_{1}}^{K}\frac{1}{k^{1/3}})\min_{K_{1}\leq k\leq K}\|\beta_{k}v_{k}\|^{2}.$ (30)
 Note that
+$\sum_{k=K_{1}}^{K}\frac{1}{k^{1/3}}\geq\int_{K_{1}}^{K+1}\frac{1}{x^{1/3}}dx=\frac{3}{2}((K+1)^{2/3}-K_{1}^{2/3}),$
+let $K=\lceil{2^{3/2}K_{1}}\rceil$ , then we obtain
+$\sum_{k=K_{1}}^{\lceil{2^{3/2}K_{1}}\rceil}\frac{1}{k^{1/3}}\geq\frac{3}{2}K_{1}^{2/3}.$ (31)
+Therefore, combining (29), (30), (31) and the bounds $Q_{K}(x_{K_{1}})\leq\overline{Q}$ and $Q_{\lceil{2^{3/2}K_{1}}\rceil+1}(x_{\lceil{2^{3/2}K_{1}}\rceil+1})\geq\underline{Q}$, we have
+$\min_{K_{1}\leq k\leq\lceil{2^{3/2}K_{1}}\rceil}\|\beta_{k}v_{k}\|^{2}\leq\frac{2l_{\max}}{3c_{1}\eta^{J_{1}}\gamma_{0}K_{1}^{2/3}}(\overline{Q}-\underline{Q}+\frac{\gamma_{0}}{2}\sigma_{y}^{2}+2(\overline{T}+1)S_{\rho}\sigma_{y}^{2}).$ (32)
+In view of definition (21) and the associated first-order optimality condition, for $\overline{y}_{k}(x_{k,0})$, we have that
+$\rho_{k}(\overline{y}_{k}(x_{k,0})-y_{k})+\gamma_{k}\overline{y}_{k}(x_{k,0})$
+$\in$ $\nabla_{y}f(x_{k,0},\overline{y}_{k}(x_{k,0}))-\partial g(\overline{y}_{k}(x_{k,0}))-\partial\delta_{S}(\overline{y}_{k}(x_{k,0})).$
+Invoking this inclusion, we further obtain that
+$\mathrm{dist}(0,\nabla_{y}f(x_{k,0},\overline{y}_{k}(x_{k,0}))-\partial g(\overline{y}_{k}(x_{k,0}))-\partial\delta_{S}(\overline{y}_{k}(x_{k,0})))$ (33)
+$\leq$ $\|\rho_{k}(\overline{y}_{k}(x_{k,0})-y_{k})+\gamma_{k}\overline{y}_{k}(x_{k,0})\|$
+$\leq$ $\sigma_{y}(2\rho_{k}+\gamma_{k}).$
+Finally, by the definition of $\rho_{k}$, $\gamma_{k}$ and $\mathcal{G}^{\beta_{k}}$ , combining (32) and (33), we conclude that
+$\min_{K_{1}\leq k\leq\lceil 2^{3/2}K_{1}\rceil}\mathcal{G}^{\beta_{k}}(x_{k,0},\overline{y}_{k}(x_{k,0}))\leq\frac{\max\{(2\rho_{0}+\gamma_{0})\sigma_{y},C\}}{K_{1}^{1/3}}$
+with
+$C=\frac{\sqrt{2l_{\max}(\overline{Q}-\underline{Q}+\frac{\gamma_{0}}{2}\sigma_{y}^{2}+2(\overline{T}+1)S_{\rho}\sigma_{y}^{2})}}{\sqrt{3c_{1}\eta^{J_{1}}\gamma_{0}}},$
+and the proof is finished. ∎
+Theorem 4.1 demonstrates that the proposed MPGDA-PA algorithm can find an $\varepsilon$-game-stationary point for problem (1) within $\mathcal{O}(\varepsilon^{-3})$ outer iterations.
+In view of Proposition 4, this also indicates that the outer iteration complexity of the MPGDA-PA algorithm for returning an $\varepsilon$-optimization-stationary point is $\mathcal{O}(\varepsilon^{-3})$.
+5 The Proposed MPGDA-PGA Algorithm
+In this section, we propose the MPGDA-PGA algorithm and establish its iteration complexity.
+In contrast to the MPGDA-PA algorithm, the MPGDA-PGA algorithm performs a single proximal gradient ascent step to update $y$, which significantly reduces the computational cost of each iteration.
+Motivated in part by xuUnifiedSingleloopAlternating2023, we develop the MPGDA-PGA algorithm by alternatively performing one manifold proximal gradient descent step and a proximal gradient ascent step for a regularized version of the original function, i.e.,
+$F_{k}(x,y):=f(x,y)+h(x)-g(y)-\frac{\gamma_{k-1}}{2}\|y\|^{2}$
+where $\gamma_{k-1}>0$ is the regularization parameter.
+More specifically, for a given pair $(x_{k},y_{k})\in\mathcal{M}\times S$ at the $k$-th iteration, the proposed MPGDA-PGA first computes a manifold proximal gradient direction $v_{k}$ similarly as in MPGDA-PA, that is:
+$v_{k}:=\mathop{\mathrm{argmin}}_{v\in T_{x_{k}}\mathcal{M}}\langle\nabla_{x}f(x_{k},y_{k}),v\rangle+h(x_{k}+v)+\frac{\beta_{k}}{2}\|v\|^{2},$ (34)
+where $\beta_{k}>0$ is the proximal parameter.
+Then, it generates $x_{k+1}:=R_{x_{k}}(\alpha_{k}v_{k})$ for some appropriate $\alpha_{k}\in(0,1]$, and accordingly set $y_{k+1}:=\hat{y}_{k}(x_{k+1})$, where the mapping $\hat{y}_{k}:\mathbb{E}_{1}\to\mathbb{E}_{2}$ is defined by
+$\hat{y}_{k}(x):=\mathop{\mathrm{argmax}}_{y\in S}\left\langle\nabla_{y}f(x,y_{k})-\gamma_{k}y_{k},y-y_{k}\right\rangle-\frac{1}{2\rho}\|y-y_{k}\|^{2}-g(y)$ (35)
+with $\rho>0$ being the proximal parameter.
+Before we propose the MPGDA-PGA algorithm formally, we establish the following lemma regarding the difference of $F_{k+1}(x_{k+1},y_{k+1})$ and $F_{k}(x_{k},y_{k})$.
+Lemma 2.
+Let $x_{k}\in\mathcal{M},~y_{k-1}\in S$ be given.
+Define $v_{k}$ as in (34), and set $x_{k+1}:=R_{x_{k}}(\alpha_{k}v_{k})$ for some $\alpha_{k}\in(0,1]$.
+Further, let $y_{k}:=\hat{y}_{k-1}(x_{k})$ and $y_{k+1}:=\hat{y}_{k}(x_{k+1})$ with $\hat{y}_{k}$ defined in (35).
+If $\rho$ and $\gamma_{k}$ satisfy
+$0<\rho\leq\frac{1-2(\kappa+1)^{-1/4}}{L_{y}},~~\gamma_{k}=\frac{2}{\rho(k+\kappa+2)^{1/4}}\text{ with }\kappa>15,$ (36)
+then it holds that
+$F_{k+1}(x_{k+1},y_{k+1})-F_{k}(x_{k},y_{k})$ (37)
+$\leq$ $-(\beta_{k}-b_{1}\alpha_{k})\alpha_{k}\|v_{k}\|^{2}+\frac{1}{2\rho}\|y_{k}-y_{k-1}\|^{2}+\frac{1}{\rho}\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}-\gamma_{k}}{2}\sigma_{y}^{2},$
+where
+$b_{1}:=M_{2}(L_{h}+G)+M_{1}^{2}(L_{x}+\rho L_{y}^{2})/2$
+with $G=\sup\{\|\nabla_{x}f(x,y)\|\mid x\in\mathcal{M},y\in S\}<\infty$, $M_{1}$ and $M_{2}>0$ being defined in Proposition 1.
+Proof.
+Let $f_{k}(x,y):=f(x,y)-\frac{\gamma_{k-1}}{2}\|y\|^{2}$. From the definition of $y_{k}$, it holds that
+$0\in\nabla_{y}f_{k}(x_{k},y_{k-1})-\partial g(y_{k})-\partial\delta_{S}(y_{k})-\frac{1}{\rho}(y_{k}-y_{k-1}).$ (38)
+By the concavity of $f_{k+1}(x,\cdot)-g(\cdot)$, it follows that
+$f_{k+1}(x_{k+1},y_{k+1})-g(y_{k+1})-(f_{k+1}(x_{k+1},y_{k})-g(y_{k}))$ (39)
+$\leq$ $\langle\nabla_{y}f_{k+1}(x_{k+1},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1})+\frac{1}{\rho}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle.$
+Note that
+$\langle\nabla_{y}f_{k+1}(x_{k+1},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle.$ (40)
+$=$ $\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
+$+\langle\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle$
+$+\langle\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1}),w_{k}\rangle$
+$-(\gamma_{k}-\gamma_{k-1})\langle y_{k},y_{k+1}-y_{k}\rangle.$
+where $w_{k}:=y_{k+1}-2y_{k}+y_{k-1}$.
+Next, we provide bounds on the inner product terms of (40).
+Firstly, invoking (Beck2017First, Theorem 5.8), by the concavity of $f_{k}(x,\cdot)$ and the Lipschitz continuity of $\nabla_{y}f_{k}(x,\cdot)$, it holds that
+$\langle\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle$ (41)
+$\leq$ $-\frac{1}{L_{y}+\gamma_{k-1}}\|\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1})\|^{2}$
+$\leq$ $-\frac{1}{L_{y}+\gamma_{-1}}\|\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1})\|^{2},$
+the last inequality comes from the fact that $\gamma_{k-1}\leq\gamma_{-1}$. In addition, by the strong concavity of $f_{k}(x,\cdot)$, we have
+$\langle\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle\leq-\gamma_{k-1}\|y_{k}-y_{k-1}\|^{2}.$ (42)
+Combining (41) and (42), it follows that
+$\langle\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle$ (43)
+$\leq$ $-\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}-\frac{1}{2(L_{y}+\gamma_{-1})}\|\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1})\|^{2}.$
+Secondly, by Cauchy-Schwarz inequality and the Lipschitz continuity of $\nabla_{y}f(x,y)$, it holds that
+$\langle\nabla_{y}f(x_{k+1},y_{k})-\nabla_{y}f(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$ (44)
+$\leq$ $\frac{\rho L_{y}^{2}}{2}\|x_{k+1}-x_{k}\|^{2}+\frac{1}{2\rho}\|y_{k+1}-y_{k}\|^{2},$
+$\langle\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1}),w_{k}\rangle$ (45)
+$\leq$ $\frac{\rho}{2}\|\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1})\|^{2}+\frac{1}{2\rho}\|w_{k}\|^{2}.$
+By the range of $\rho$ and direct calculation, it holds that $\rho<\frac{1}{L_{y}+\gamma_{-1}}$.
+Combining this, the definition of $f_{k}(x_{k+1},y_{k+1})$, (39), (40), (43), (44), (45), and the following identities
+$\frac{1}{\rho}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle$ $=\frac{1}{2\rho}(\|y_{k+1}-y_{k}\|^{2}+\|y_{k}-y_{k-1}\|^{2}-\|w_{k}\|^{2}),$ (46)
+$(\gamma_{k}-\gamma_{k-1})\langle y_{k},y_{k+1}-y_{k}\rangle$ $=\frac{\gamma_{k}-\gamma_{k-1}}{2}(\|y_{k+1}\|^{2}-\|y_{k}\|^{2}-\|y_{k+1}-y_{k}\|^{2}),$ (47)
+we deduce that
+$f(x_{k+1},y_{k+1})-g(y_{k+1})-(f(x_{k+1},y_{k})-g(y_{k}))$ (48)
+$\leq$ $\frac{\rho L_{y}^{2}}{2}\|x_{k+1}-x_{k}\|^{2}+\frac{1}{\rho}\|y_{k+1}-y_{k}\|^{2}+\frac{1}{2\rho}\|y_{k}-y_{k-1}\|^{2}$
+$+\frac{\gamma_{k-1}}{2}(\|y_{k+1}\|^{2}-\|y_{k}\|^{2}).$
+On the other hand, following an argument similar to the proof of (chenProximalGradientMethod2020, Lemma 5.2), we have
+$f(x_{k+1},y_{k})+h(x_{k+1})\leq f(x_{k},y_{k})+h(x_{k})-(\frac{\beta_{k}}{\alpha_{k}}-b_{2})\|\alpha_{k}v_{k}\|^{2},$
+where $b_{2}=M_{2}(L_{h}+G)+M_{1}^{2}L_{x}/2$. By Proposition 1, this and inequality (48) yield inequality (37). ∎
+However, Lemma 2 alone does not yield an upper bound on a positively weighted sum of $\|v_{k}\|^{2}$ and $\|y_{k+1}-y_{k}\|^{2}$, which provide an upper bound for $\mathcal{G}^{\beta_{k}}(x_{k},y_{k})$ defined in (13).
+To address this difficulty, we further refine the result in (37).
+Specifically, in the next proposition, we derive a new inequality (see (52)) to better elucidate the relationship between $\|v_{k}\|^{2}$ and $\|y_{k+1}-y_{k}\|^{2}$.
+Furthermore, by using this new inequality, we construct a new value function (see (50)) which involves $F_{k}$ and some additional terms associated with $\|y-y_{k-1}\|^{2}$ and $\|y\|^{2}$.
+Proposition 7.
+Let $x_{k}\in\mathcal{M},~y_{k-1}\in S$ be given.
+Define $v_{k}$ as in (34), and set $x_{k+1}:=R_{x_{k}}(\alpha_{k}v_{k})$ for some $\alpha_{k}\in(0,1]$.
+Further, let $y_{k}:=\hat{y}_{k-1}(x_{k})$ and $y_{k+1}:=\hat{y}_{k}(x_{k+1})$ with $\hat{y}_{k}$ defined in (35).
+Also denote
+$\mathcal{H}_{k}(y)$ $:=\left(\frac{4}{\rho^{2}\gamma_{k}}-\frac{4}{\rho}\right)\|y-y_{k-1}\|^{2}+\frac{4}{\rho}\left(1-\frac{\gamma_{k-1}}{\gamma_{k}}\right)\|y\|^{2},$ (49)
+${\mathcal{F}}_{k}(x,y)$ $:=F_{k}(x,y)+\frac{1}{2\rho}\|y-y_{k-1}\|^{2}+(\frac{4}{\rho}\frac{\gamma_{k-1}}{\gamma_{k}}+\frac{\gamma_{k-1}}{2})\sigma_{y}^{2}+\mathcal{H}_{k}(y).$ (50)
+If $\rho$ and $\gamma_{k}$ satisfy (36), then it holds that
+$\mathcal{F}_{k+1}(x_{k+1},y_{k+1})-\mathcal{F}_{k}(x_{k},y_{k})$ (51)
+$\leq$ $-\big(\beta_{k}-(b_{1}+\frac{8L_{y}^{2}M_{1}^{2}}{\rho\gamma_{k}^{2}})\alpha_{k}\big)\alpha_{k}\|v_{k}\|^{2}-\frac{1}{10\rho}\|y_{k+1}-y_{k}\|^{2}.$
+Proof.
+Let $f_{k}(x,y):=f(x,y)-\frac{\gamma_{k-1}}{2}\|y\|^{2}$. By (38) and the concavity of g-g, we have
+$\frac{1}{\rho}\langle w_{k},y_{k+1}-y_{k}\rangle\leq$ $\langle\nabla_{y}f_{k+1}(x_{k+1},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle,$
+where $w_{k}=y_{k+1}-2y_{k}+y_{k-1}$.
+Similar to the proof in Lemma 2, combining this with (40), (43), (44) (replacing $\rho$ by $2/\gamma_{k}$), (45), (47), $\gamma_{k}\leq\gamma_{k-1}$, and the following identity
+$\frac{1}{\rho}\langle w_{k},y_{k+1}-y_{k}\rangle=\frac{1}{2\rho}(\|y_{k+1}-y_{k}\|^{2}-\|y_{k}-y_{k-1}\|^{2}+\|w_{k}\|^{2}),$
+we deduce that
+$\frac{1}{2\rho}\|y_{k+1}-y_{k}\|^{2}\leq$ $\frac{1}{2\rho}\|y_{k}-y_{k-1}\|^{2}+\frac{\gamma_{k}}{4}\|y_{k+1}-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|y_{k}-y_{k-1}\|^{2}$ (52)
+$+\frac{L_{y}^{2}}{\gamma_{k}}\|x_{k+1}-x_{k}\|^{2}+\frac{\gamma_{k-1}-\gamma_{k}}{2}(\|y_{k+1}\|^{2}-\|y_{k}\|^{2})$
+$\leq$ $\frac{1}{2\rho}\|y_{k}-y_{k-1}\|^{2}+\frac{\gamma_{k}}{4}\|y_{k+1}-y_{k}\|^{2}-\frac{\gamma_{k}}{2}\|y_{k}-y_{k-1}\|^{2}$
+$+\frac{\alpha_{k}^{2}M_{1}^{2}L_{y}^{2}}{\gamma_{k}}\|v_{k}\|^{2}+\frac{\gamma_{k-1}-\gamma_{k}}{2}(\|y_{k+1}\|^{2}-\|y_{k}\|^{2}),$
+where the last inequality comes from Proposition 1. Multiplying (52) by $\frac{8}{\rho\gamma_{k}}$ and then rearranging it appropriately, we obtain
+$\mathcal{H}_{k+1}(y_{k+1})\leq$ $\mathcal{H}_{k}(y_{k})+\frac{8\alpha_{k}^{2}M_{1}^{2}L_{y}^{2}}{\rho\gamma_{k}^{2}}\|v_{k}\|^{2}+\frac{4}{\rho}(\frac{\gamma_{k-1}}{\gamma_{k}}-\frac{\gamma_{k}}{\gamma_{k+1}})\|y_{k+1}\|^{2}$ (53)
+$-\big(\frac{2}{\rho}-\frac{4}{\rho^{2}}(\frac{1}{\gamma_{k+1}}-\frac{1}{\gamma_{k}})\big)\|y_{k+1}-y_{k}\|^{2}.$
+Since $\frac{\gamma_{k-1}}{\gamma_{k}}>\frac{\gamma_{k}}{\gamma_{k+1}}$ and for $\kappa>15$,
+$\frac{1}{\rho}(\frac{1}{\gamma_{k}}-\frac{1}{\gamma_{k-1}})=\frac{1}{2}\big((k+\kappa+2)^{1/4}-(k+\kappa+1)^{1/4}\big)<\frac{1}{2}\big(3^{1/4}-2^{1/4}\big)<\frac{1}{10},$
+by (37) and (53), we obtain (51). ∎
+Proposition 7 implies that we can ensure the sufficient descent of $\mathcal{F}_{k}(x_{k},y_{k})$ by choosing a small enough stepsize $\alpha_{k}\in(0,1]$.
+For possible acceleration, we further incorporate a backtracking line search procedure regarding inequality (51) to determine $\alpha_{k}$.
+The MPGDA-PGA algorithm is formally presented in Algorithm 2.
+Algorithm 2 Manifold Proximal Gradient Descent Ascent-Proximal Gradient Ascent (MPGDA-PGA)
 […]
-let $K=\lceil{2^{3/2}K_{1}}\rceil$ , then we obtain
-$\sum_{K_{1}\leq k\leq\lceil{2^{3/2}K_{1}}\rceil}\frac{1}{k^{1/3}}\geq\frac{3}{2}K_{1}^{2/3}.$ (22)
-Therefore, combining (20), (21) and (22), with $Q_{K}(x_{K_{1}})\leq\overline{Q}$ and $Q_{\lceil{2^{3/2}K_{1}}\rceil}(x_{\lceil{2^{3/2}K_{1}}\rceil})\geq\underline{Q}$, we have
-$\min_{K_{1}\leq k\leq\lceil{2^{3/2}K_{1}}\rceil}\|\beta_{k}v_{k}\|^{2}\leq\frac{2l_{\max}}{3\bar{\alpha}\eta\gamma_{0}K_{1}^{2/3}}(\overline{Q}-\underline{Q}+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2}).$ (23)
-In view of definition (15) and the associated first-order optimality condition, for $\bar{y}_{k}(x_{k,0})$, we have that
-$\rho_{k}(\bar{y}_{k}(x_{k,0})-y_{k})+\gamma_{k}\bar{y}_{k}(x_{k,0})\in\nabla_{y}f(x_{k,0},\bar{y}_{k}(x_{k,0}))-\partial g(\bar{y}_{k}(x_{k,0}))-\partial\delta_{S}(\bar{y}_{k}(x_{k,0})).$
-Invoking this inclusion, we further obtain that
-$\mathrm{dist}(0,\nabla_{y}f(x_{k,0},\bar{y}_{k}(x_{k,0}))-\partial g(\bar{y}_{k}(x_{k,0}))-\partial\delta_{S}(\bar{y}_{k}(x_{k,0})))$ (24a)
-$\leq$ $\|\rho_{k}(\bar{y}_{k}(x_{k,0})-y_{k})+\gamma_{k}\bar{y}_{k}(x_{k,0})\|$ (24b)
-$\leq$ $\sigma_{y}(2\rho_{k}+\gamma_{k}).$ (24c)
-Finally, by the definition of $\rho_{k}$, $\gamma_{k}$ and $\mathcal{G}^{\beta_{k}}$ , combining (23) and (24), we conclude that
-$\min_{K_{1}\leq k\leq\lceil 2^{3/2}K_{1}\rceil}\mathcal{G}^{\beta_{k}}(x_{k,0},\bar{y}_{k}(x_{k,0}))\leq\frac{\max\{(2\rho_{0}+\gamma_{0})\sigma_{y},C\}}{K_{1}^{1/3}}$
-with
-$C=\frac{\sqrt{2(\overline{Q}-\underline{Q}+2(\overline{T}+2)S_{\rho}\sigma_{y}^{2})l_{\max}}}{\sqrt{3\bar{\alpha}\eta\gamma_{0}}},$
+1: Set $y_{0}=\hat{y}_{-1}(x_{0})$.
+2:for $k=0,1,\dots$ do
+3:  Choose $l_{k}\in[\hat{l}_{\min},\hat{l}_{\max}]$ and set $\beta_{k}=l_{k}/\gamma_{k}^{2}$.
+4:  Compute $v_{k}$ by (34).
+5:  Find the smallest non-negative integer $j$ such that $x_{k}^{j}:=R_{x_{k}}(\eta^{j}v_{k})$ satisfies
+${\mathcal{F}}_{k+1}(x_{k}^{j},\hat{y}_{k}(x_{k}^{j}))\leq{\mathcal{F}}_{k}(x_{k},y_{k})-c_{1}\eta^{j}\beta_{k}\|v_{k}\|^{2}-\frac{1}{10\rho}\|\hat{y}_{k}(x_{k}^{j})-y_{k}\|^{2},$ (54)
+6:  Set $x_{k+1}=x_{k}^{j},~y_{k+1}=\hat{y}_{k}(x_{k+1})$.
+7:end for
+8:return $(x_{k+1},y_{k+1})$
+Before proceeding, we make some remarks on the iteration procedure of the MPGDA-PGA algorithm. First, the update of $y_{k+1}$ in MPGDA-PGA can be formulated as:
+$y_{k+1}=\mathop{\mathrm{argmin}}_{y\in S}\left\{g(y)+\frac{1}{2\rho}\|y-\left((1-\rho\gamma_{k})y_{k}+\rho\nabla_{y}f(x_{k+1},y_{k})\right)\|^{2}\right\},$ (55)
+which is a proximal problem of the type (2b) and can be evaluated exactly and efficiently.
+Second, in practice, we use some special choices of $l_{k}\in[\hat{l}_{\min},\hat{l}_{\max}]$ to enhance the efficiency of MPGDA-PGA in the experiments.
+Motivated by the Riemannian Barzilai-Borwein (BB) stepsize iannazzo2018riemannian; 2013AWen, we choose $l_{k}$ as follows:
+$l_{k}=\begin{cases}\min\left\{\max\left\{\hat{l}_{\min},\gamma_{k}^{2}\left|\frac{\langle\Delta X_{k},\Delta R_{k}\rangle}{\|\Delta X_{k}\|^{2}}\right|\right\},\hat{l}_{\max}\right\},&\text{if }\Delta X_{k}\neq 0,\\ \hat{l}_{\max},&\text{if }\Delta X_{k}=0,\end{cases}$
+where $\Delta X_{k}=x_{k}-x_{k-1}$ and $\Delta R_{k}=\mathrm{grad}f(x_{k},y_{k})-\mathrm{grad}f(x_{k-1},y_{k-1})$.
+By Proposition 7, the backtracking line search procedure in Algorithm 2 is well defined; i.e., it will terminate after a finite number of steps. We present this result in the following proposition.
+Proposition 8.
+For all $k\geq 0$, the condition (54) in Algorithm 2 is satisfied within at most $J_{2}$ backtracking steps, where
+$J_{2}:=\max\left\{\left\lceil\log_{\eta}\left(\frac{\rho(1-c_{1})\hat{l}_{\min}}{\rho b_{1}\gamma_{-1}^{2}+8L_{y}^{2}M_{1}^{2}}\right)\right\rceil,0\right\}$
+with $b_{1}$ being defined in Lemma 2.
+Below we establish the iteration complexity for Algorithm 2. To this end, we first define two constants $\overline{\mathcal{F}}$ and $\underline{\mathcal{F}}$ as
+$\overline{\mathcal{F}}$ $:=\sup\{{F(x,y)+(\frac{6}{\rho}+\frac{16}{\rho^{2}\gamma_{0}}+\frac{4\gamma_{-1}}{\rho\gamma_{0}}+\frac{\gamma_{-1}}{2})\sigma_{y}^{2}\mid x\in\mathcal{M},y\in S}\},$
+$\underline{\mathcal{F}}$ $:=\inf\{F(x,y)-(\frac{\gamma_{-1}}{2}+\frac{24}{\rho})\sigma_{y}^{2}\mid x\in\mathcal{M},y\in S\}.$
+By direct calculation, we have $\underline{\mathcal{F}}\leq\mathcal{F}_{k}(x,y)$ for all $k\geq 0$ and $\mathcal{F}_{0}(x,y)\leq\overline{\mathcal{F}}$ for any point $(x,y)\in\mathcal{M}\times(S)$.
+Theorem 5.1.
+Given $\varepsilon>0$, let $\{(x_{k},y_{k})\}$ be generated by Algorithm 2 and $T(\varepsilon)$ be defined in (28). Then, it holds that
+$T(\varepsilon)\leq\frac{\big((15+4\kappa)^{1/4}(2(\overline{\mathcal{F}}-\underline{\mathcal{F}})/C_{1})^{1/2}(\rho L_{y}+1)+2\sigma_{y}/\rho\big)^{4}}{\varepsilon^{4}},$
+where
+$C_{1}=\min\{\frac{4c_{1}\eta^{J_{2}}}{\hat{l}_{\max}\rho^{2}},\frac{\rho(\kappa+2)^{1/2}}{10}\}.$
+Here, the constants $c_{1},~\eta,~\kappa,~\hat{l}_{\max}$, and $\rho$ are parameters of Algorithm 2, and $J_{2}$ is defined in Proposition 8.
+Proof.
+Summing the inequality (54) over $k=0,1,\dots,K$, it holds that
+$\sum_{k=0}^{K}c_{1}\eta^{J_{2}}\beta_{k}\|v_{k}\|^{2}+\frac{1}{10\rho}\|y_{k+1}-y_{k}\|^{2}\leq\overline{\mathcal{F}}-\underline{\mathcal{F}},$ (56)
+which follows from Proposition 8 and the definition of $\overline{\mathcal{F}}$ and $\underline{\mathcal{F}}$. Combining (56) and the definition of $\beta_{k}$, we have
+$(\sum_{k=0}^{K}\frac{C_{1}}{(k+\kappa+2)^{1/2}})\min_{0\leq k\leq K}(\|\beta_{k}v_{k}\|^{2}+\|\frac{1}{\rho}(y_{k}-y_{k-1})\|^{2})$ $\leq\overline{\mathcal{F}}-\underline$$\mathcal{F}$}, (57)
+where $C_{1}=\min\{\frac{4c_{1}\eta^{J_{2}}}{\hat{l}_{\max}\rho^{2}},\frac{\rho(\kappa+2)^{1/2}}{10}\}$. Note that
+$\sum_{k=0}^{K}\frac{1}{(k+\kappa+2)^{1/2}}$ $\geq\int_{1}^{K+1}\frac{1}{(x+\kappa+2)^{1/2}}dx$
+$=\frac{1}{2}((K+\kappa+3)^{1/2}-(\kappa+3)^{1/2}),$
+this and inequality (57) yield that
+$\min_{0\leq k\leq K}\|\beta_{k}v_{k}\|^{2}+\|\frac{1}{\rho}(y_{k}-y_{k-1})\|^{2}$ $\leq\frac{2(\overline{\mathcal{F}}-\underline{\mathcal{F}})}{C_{1}\big((K+\kappa+3)^{1/2}-(\kappa+3)^{1/2}\big)}.$ (58)
+On the other hand, by the optimal condition of definition of $y_{k}$, for $k\geq 2$, we have
+$\mathrm{dist}(0,\nabla_{y}f(x_{k},y_{k})-\partial g(y_{k})-\partial\delta_{S}(y_{k}))\leq(L_{y}+\frac{1}{\rho})\|y_{k}-y_{k-1}\|+\gamma_{k-1}\|y_{k}\|.$
+Combining this bound with the definition of $\gamma_{k-1}$ and inequality (58), we obtain that
+$\min_{0\leq k\leq K}\mathcal{G}^{\beta_{k}}(x_{k},y_{k})\leq\frac{(2(\overline{\mathcal{F}}-\underline{\mathcal{F}})/C_{1})^{1/2}(\rho L_{y}+1)}{((K+\kappa+3)^{1/2}-(\kappa+3)^{1/2})^{1/2}}+\frac{2\sigma_{y}}{\rho(K+2)^{1/4}}.$ (59)
+By the fact that $(K+\kappa+3)^{1/2}-(\kappa+3)^{1/2}\geq\frac{K^{1/2}}{(15+4\kappa)^{1/2}}$ for $\kappa>15$ and any $K\geq 1$, we deduce from (59) that
+$\min_{0\leq k\leq K}\mathcal{G}^{\beta_{k}}(x_{k},y_{k})\leq\frac{(15+4\kappa)^{1/4}(2(\overline{\mathcal{F}}-\underline{\mathcal{F}})/C_{1})^{1/2}(\rho L_{y}+1)+2\sigma_{y}/\rho}{K^{1/4}},$
 and the proof is finished. ∎
-Theorem 4.1 demonstrates that the proposed MPGDA algorithm can find an $\varepsilon$-game-stationary point for problem (1) within $O(\varepsilon^{-3})$ iterations.
-In view of Proposition 4, this also indicates that the iteration complexity of the MPGDA algorithm for returning an $\varepsilon$-optimization-stationary point is $\mathcal{O}(\varepsilon^{-3})$.
+Theorem 5.1 demonstrates that the proposed MPGDA-PGA algorithm can find an $\varepsilon$-game-stationary point for problem (1) within $\mathcal{O}(\varepsilon^{-4})$ outer iterations.
+In view of Proposition 4, this also indicates that the outer iteration complexity of the MPGDA-PGA algorithm for returning an $\varepsilon$-optimization-stationary point is $\mathcal{O}(\varepsilon^{-4})$.
 6 Numerical experiments
-In this section, we demonstrate the efficiency of the proposed MPGDA algorithm by applying it to FSPCA problem (4) and SSC problem (7).
-All the experiments are implemented in MATLAB R2019a and conducted on a standard PC with 3.40GHz Intel(R) Core(TM) i5-7500 CPU and 16GB of RAM.
-We first specify some implementation details of the proposed algorithm.
-Throughout the tests, we set $c_{1}=10^{-4}$, $\eta=0.1$, $l_{\min}=10^{-16}$, $l_{\max}=10^{16}$, $\tau_{1}=0.999$ and $\tau_{2}=0.9$.
-The parameters $\gamma_{0}$, $\xi_{0}>0$, $\theta>1$ and $\{T_{k}\}$ vary across the two application problems.
-Since $h$ is the $\ell_{1}$-norm in both problems, the semismooth Newton method developed in (chenProximalGradientMethod2020, , Section 4.2) is used to solve subproblem (17).
+In this section, we illustrate the proposed algorithms via three numerical examples.
+We begin with an analytic example of problem (1), where the objective function is not linear with respect to $y$.
+This example is used to illustrate the differences in behavior between the MPGDA-PA algorithm and the MPGDA-PGA algorithm.
+Then, we examine the performance of the MPGDA-PA algorithm for solving the FSPCA problem (5) and the SSC problem (8) respectively.
+Since these two problems are nonconvex-linear, the two proposed algorithms have comparable computational cost at each iteration, but the MPGDA-PA algorithm enjoys better outer iteration complexity.
+Hence, it is more reasonable to apply MPGDA-PA rather than MPGDA-PGA algorithm for solving problems (5) and (8).
+All experiments are implemented in MATLAB R2025a and conducted on a standard PC with 3.40GHz Intel(R) Core(TM) i5-7500 CPU and 24GB of RAM.
+We first specify some implementation details of the proposed MPGDA-PA algorithm.
+Throughout the tests, we set $c_{1}=10^{-4}$, $\eta=0.1$, $l_{\min}=10^{-16}$, $l_{\max}=10^{16}$, $\delta_{0}=10^{10}$, $\tau_{1}=0.999$, and $\tau_{2}=0.9$.
+The parameters $\gamma_{0}$, $\xi_{0}>0$, $\theta>1$ and $\{T_{k}\}$ vary across the three examples.
+Since $h$ is the $\ell_{1}$-norm in both problems, the semismooth Newton method developed in (chenProximalGradientMethod2020, Section 4.2) is used to solve subproblem (23).
 We utilize the QR factorization as the retraction on the Stiefel manifold 2008Optimization.
+6.1 Experiments on an Analytic Minimax Problem
+Consider the following nonconvex-concave minimax problem on $\mathbb{R}^{2}\times\mathbb{R}$:
+$\min_{x_{1}^{2}+x_{2}^{2}=1}\max_{0.3\leq y\leq 1}-0.01x_{1}^{3}y-y\ln(y),$ (60)
+which is a special case of problem (1) with $\mathcal{M}$ being the Stiefel manifold $\mathrm{St}(2,1)$, $S$ being the interval $[0.3,1]$, and $f(x,y)=-0.01x_{1}^{3}y-y\ln(y)$.
+A direct verification shows that $(1,0)$ is the unique minimizer to the value function of (60) and $(x^{*},y^{*}):=(1,0,e^{-1.01})$ is a game-stationary point.
+Now we apply the two proposed algorithms to solve (60).
+For the MPGDA-PA algorithm, we set $\gamma_{0}=0.005$, $\xi_{0}=1$, $\theta=1.5$, and $T_{k}=1$.
+We use the bisection method to solve the corresponding $y$-subproblem (21), which is implemented via the fminbnd command in MATLAB.
+For MPGDA-PGA, we set $c_{1}=10^{-4}$, $\hat{l}_{\min}=10^{-16}$, $\hat{l}_{\max}=10^{8}$, $\rho=0.2$, $\kappa=10^{16}$, and $\eta=0.5$.
+Both algorithms are initialized from the same starting point $(0.8,0.6,0.3)\in\mathrm{St}(2,1)\times[0.3,1]$.
+We use the Euclidean distance between the iterates $(x_{k},y_{k})$ and the optimal point $(x^{*},y^{*})$, denoted by $D_{k}$, to measure the accuracy of the iterates generated by them.
+Table 1: Comparison of iteration numbers and CPU time for reaching different levels of distance
+Level Iter. Time
+MPGDA-PA MPGDA-PGA MPGDA-PA MPGDA-PGA
+$1\times 10^{-2}$ 17 918 0.038659 0.023850
+$1\times 10^{-3}$ 19 2100 0.044875 0.054841
+$3\times 10^{-4}$ 21 2767 0.058287 0.073542
+$2\times 10^{-4}$ 38 3067 0.114491 0.084889
+$1.5\times 10^{-4}$ 88 3455 0.167304 0.118553
+Table 1 presents the first outer iteration number ( Iter.) required by the proposed algorithms to reach different levels of distance, and the corresponding CPU time in seconds (Time).
+Moreover, Figures (1(a)) and (1(b)) show $D_{k}$ (in logarithmic scale) versus iteration number $k$ and $D_{k}$ (in logarithmic scale) versus CPU time in seconds respectively.
+We observe that the iterates generated by both proposed algorithms converge to the optimal solution $(x^{*},y^{*})$.
+Although the MPGDA-PA algorithm requires to solve a nonlinear and strongly concave subproblem with respect to $y$ in each iteration, which results in a higher computational cost per iteration, it converges in substantially fewer outer iterations.
+In contrast, MPGDA-PGA performs simpler iterations at a much lower cost per iteration, but requires much more iterations to achieve comparable accuracy.
+Consequently, the overall CPU times of both algorithms are similar across different levels of accuracy, demonstrating a consistent trade-off between per-iteration cost and convergence speed for solving problem (60).
+(a) $D_{k}$ (in log scale) vs Iters
+(b) $D_{k}$ (in log scale) vs Time
+Figure 1: Plots of $D_{k}$ (in log scale) for the proposed algorithms.
 6.2 Experiments on FSPCA
-In this subsection, we compare the proposed MPGDA algorithm with the constraint relaxed majorization-minimization (CRMM) algorithm (babuFairPrincipalComponent2023, , Algorithm 3) for solving FSPCA problem (4).
+In this subsection, we compare the proposed MPGDA-PA algorithm with the constraint relaxed majorization-minimization (CRMM) algorithm (babuFairPrincipalComponent2023, Algorithm 3) for solving FSPCA problem (5).
 The CRMM algorithm first relaxes the semi-orthogonality constraint $X^{T}X=I$ into a linear matrix inequality constraint and then applies some majorization-minimization iteration procedure, which requires solving a quadratic program at each iteration.
 We perform tests on synthetic datasets and a real dataset where the samples belong to two groups.
-The synthetic datasets are generated following similar settings to (2016SparseGeneralizedEigenvalueProblem, , Section 5.1).
-In detail, for the $i$-th group, we randomly generate $m_{i}=200$ samples with $d=40$ features according to Gaussian distribution with mean $\mu^{i}\in\mathbb{R}^{d}$ and covariance $\Sigma\in\mathbb{R}^{d\times d}$.
+The synthetic datasets are generated following similar settings to (2016SparseGeneralizedEigenvalueProblem, Section 5.1).
+In detail, for the $i$-th group, we randomly generate $m_{i}=200$ samples with $d=40$ features according to the Gaussian distribution with mean $\mu^{i}\in\mathbb{R}^{d}$ and covariance $\Sigma\in\mathbb{R}^{d\times d}$.
 We set $\mu_{j}^{1}=0,\ \mu_{j}^{2}=1/3$, for $j\in\{2,4,\cdots,200\}$, and $\mu_{j}^{i}=0$ otherwise.
@@
 Furthermore, the CRMM algorithm is terminated if the relative error between the solutions obtained over two successive iterations is less than $10^{-6}$.
-For MPGDA algorithm, we set $\gamma_{0}=10^{-6}$, $\xi_{0}=4\sqrt{r}*10^{4}$, $\theta=1.5$, $T_{k}\equiv 15$, and terminate the algorithm if it satisfies:
-$\mathcal{G}^{\beta_{k,0}}(x_{k},y_{k})<\varepsilon$ (25)
-with $\varepsilon=10^{-6}$. Both compared algorithms start from the same random initial points on $\mathrm{St}(d,r)$ and the maximum iteration unmber of them are set as 1000.
+For the MPGDA-PA algorithm, we set $\gamma_{0}=10^{-6}$, $\xi_{0}=4\sqrt{r}*10^{4}$, $\theta=1.5$, $T_{k}\equiv 15$, and we terminate the algorithm if it satisfies:
+$\mathcal{G}^{\beta_{k,0}}(x_{k},y_{k})<\varepsilon\text{ with }\varepsilon=10^{-6}.$ (61)
+Both compared algorithms start from the same random initial points on $\mathrm{St}(d,r)$ and the maximum iteration number of them are set as 1000.
 We present the average results over 50 runs on different synthetic datasets and results for the default credit dataset in Table 2.
-We report the objective values (Obj), and CPU time in seconds (Time) and the number of iterations (Iter).
-It is observed that MPGDA algorithm achieves similar objective values to the CRMM algorithm while requiring much less computational time.
-Specially, our proposed algorithm is on average more than 20 times faster than the CRMM algorithm.
+We report the objective value (Obj), CPU time in seconds (Time) and the number of iterations (Iter).
+It is observed that the MPGDA-PA algorithm achieves similar objective values to the CRMM algorithm while requiring much less computational time.
+Specifically, our proposed algorithm is on average more than 20 times faster than the CRMM algorithm.
 Table 2: Results for FSPCA
-CRMM MPGDA
-Dataset r Obj. Time (s) Iter. Obj. Time (s) Iter.
-Synthetic 2 $-9.795$ $0.269$ $201$9.808 0.007 167
-3 $-14.380$ $0.344$ 155 -14.382 0.010 $192$
-4 $-18.714$ $0.443$ 162 -18.745 0.012 $217$
-5 $-22.774$ $0.209$ 83 -22.809 0.008 $197$
-Credit 2 $-9.531$ $4.014$ 93 -9.742 0.133 $120$
-3 $-10.557$ $15.692$ $271$10.802 0.104 91
-4 $-11.637$ $8.977$ $142$11.923 0.010 5
-5 $-12.572$ $20.058$ $277$12.724 0.012 4
+CRMM MPGDA-PA
+Dataset r Obj. Time Iter. Obj. Time Iter.
+Synthetic 2 9.802-9.802 $0.122$ $204$ 9.820-9.820 0.005 $167$
+3 14.399-14.399 $0.131$ $165$ 14.401-14.401 0.007 $189$
+4 18.755-18.755 $0.175$ $148$ 18.776-18.776 0.008 $216$
+5 22.840-22.840 $0.089$ $79$ 22.867-22.867 0.007 $216$
+Credit 2 9.531-9.531 $4.305$ $93$ 9.742-9.742 0.276 $124$
+3 10.557-10.557 $14.599$ $272$ 10.802-10.802 0.138 $91$
+4 11.637-11.637 $6.478$ $143$ 11.923-11.923 0.013 55
+5 12.572-12.572 $14.332$ $274$ 12.724-12.724 0.013 44
 6.3 Experiments on SSC
-In this subsection, we compare the proposed MPGDA method with the RADA-RGD and RADA-PGD algorithms in xuRiemannianAlternatingDescent2024 for solving the SSC problem.
+In this subsection, we compare the proposed MPGDA-PA method with the RADA-RGD and RADA-PGD algorithms in xuRiemannianAlternatingDescent2024 for solving the SSC problem.
 Note that in xuRiemannianAlternatingDescent2024, the original SSC problem (6) is reformulated into the following smooth minimax problem on the Grassmann manifold so that the RADA-PGD and RADA-RGD algorithms can be directly applied:
@@
 Here, $\mathrm{Gr}(N,p)=\{XX^{\top}\in\mathbb{R}^{N\times N}\mid X\in\mathrm{St}(N,p)\}$ denotes the Grassmann manifold AGrassmannmanifoldhandbook2024A.
-The codes of these two algorithms for solving (26) are downloaded from: https://github.com/XuMeng00124/RADAopt.
+The code of these two algorithms for solving (62) is downloaded from: https://github.com/XuMeng00124/RADAopt.
 In the tests, we generate synthetic datasets as in 2016ConvexSparseSpectralClustering, where the data points $\{a_{i}\}_{i=1}^{N}$ are independently drawn from the standard Gaussian distribution and $W_{ij}=|\langle a_{i},a_{j}\rangle|$.
-For the two RADA algorithms, we adopt the parameters as suggested in (xuRiemannianAlternatingDescent2024, , Section 6.3).
+For the two RADA algorithms, we adopt the parameters as suggested in (xuRiemannianAlternatingDescent2024, Section 6.3) except the number of inner iterations ($T_{k}$).
 Specifically, they are terminated once an $\varepsilon$-game-stationary point of problem (62) is found with $\varepsilon=10^{-4}$ and the initial point is set as $X_{0}X_{0}^{\top}$, where $X_{0}$ consists of $p$ eigenvectors associated with the $p$ smallest eigenvalues of $L$.
-For MPGDA algorithm, we set $\gamma_{0}=10^{-5}$, $\xi_{0}=\sqrt{p}N^{2}$, $\theta=2$ and $T_{k}\equiv 3$.
-Moreover, we choose the aforementioned $X_{0}$ as the initial point of MPGDA algorithm and terminate the method if it satisfies (25) with $\varepsilon=10^{-4}$.
+For the MPGDA-PA algorithm, we set $\gamma_{0}=10^{-5}$, $\xi_{0}=\sqrt{p}N^{2}$, and $\theta=2$.
+Moreover, we choose the aforementioned $X_{0}$ as the initial point of the MPGDA-PA algorithm and terminate the method if it satisfies (61) with $\varepsilon=10^{-4}$.
 The maximum number of iterations for all the compared algorithms is set as 1000.
+In the experiments, we generate 50 synthetic datasets with $N=200$ data points.
+To fine tune the parameters $T_{k}$ involved, we test all the compared algorithms with $T_{k}\in\{1,2,3,4,5\}$ for solving the SSC problem with various choices of $p$ and $\mu$.
+The results reveal that the RADA-PGD, RADA-RGD and MPGDA-PA algorithms consistently achieve optimal performance with $T_{k}=1$, $2$, and $3$, respectively.
+Due to the limitation of space, we only report one representative result in Table 3, which shows the performance evaluation for different choices of $T_{k}$ on the SSC problem with $(p,\mu)=(5,1.0)$.
+Also, the computational results for the compared algorithms with their respective optimal $T_{k}$ are summarized in Table 4.
+Both tables present the objective function value of problem (6) (Obj.), the CPU time in seconds (Time), and number of outer iterations (Iter.) averaged over the 50 datasets.
+We observe that all the compared algorithms attain comparable objective function values for different numbers of groups and regularization parameters $\mu$.
+In addition, the proposed MPGDA-PA method consistently outperforms the two RADA algorithms in terms of CPU time, which again validates its computational efficiency.
+The superior performance of the proposed MPGDA-PA algorithm may be attributed to the fact that it is able to directly address the nonsmooth $\ell_{1}$ norm in the SSC problem, while the RADA algorithms can only tackle the converted smooth minimax formulation of SSC.
+Table 3: Performance evaluation for different choices of $T_{k}$ on the SSC problem with $(p,\mu)=(5,1.0)$
+RADA-PGD RADA-RGD MPGDA-PA
+$T_{k}$ Obj. Time Iter. Obj. Time Iter. Obj. Time Iter.
+1 $9.971$ 0.578 $300$ $17.590$ $1.787$ $1000$ $22.320$ $2.063$ $1000$
+2 $9.971$ $1.148$ $300$ $10.070$ 0.873 $238$ $10.029$ $2.080$ $848$
+3 $9.971$ $1.629$ $297$ $9.967$ $1.170$ $217$ $9.968$ 0.256 $97$
+4 $9.971$ $2.198$ $301$ $9.967$ $1.830$ $254$ $9.968$ 0.298 $88$
+5 $9.971$ $2.516$ $300$ $64.748$ $2.742$ $289$ $9.970$ 0.311 $84$
 Table 4: Results for SSC
-Config Para $(p/\mu)$ RADA-PGD RADA-RGD MPGDA
+Parameters
+($p$, $\mu$)
+RADA-PGD RADA-RGD MPGDA-PA
 Obj. Time Iter. Obj. Time Iter. Obj. Time Iter.
-$\mu=0.1$ $N=200$ 2 2.187 $0.513$ $265$ 2.187 $0.796$ $153$ 2.187 0.114 65
-4 $4.374$ $0.530$ $259$ 4.373 $0.812$ $157$ $4.374$ 0.116 64
-6 6.560 $0.521$ $224$ 6.560 $0.855$ $164$ 6.560 0.117 63
-8 8.747 $0.580$ $198$ 8.747 $0.861$ $166$ 8.747 0.129 64
-10 10.934 $0.621$ $180$ 10.934 $0.995$ $167$ 10.934 0.129 65
-$p=5$ $N=200$ 0.1 5.467 $0.535$ $244$ 5.467 $0.868$ $164$ 5.467 0.122 63
-0.2 5.967 $0.556$ $263$ 5.967 $0.934$ $178$ 5.967 0.133 71
-0.5 7.467 $0.593$ $286$ 7.467 $1.071$ $202$ 7.467 0.174 86
-1.0 9.967 $0.616$ $300$ 9.967 $1.188$ $228$ 9.967 0.205 95
-The computational results averaged over 50 synthetic data sets with $N=200$ data points are summarized in Table 2, where we report the obtained objective function value of problem (5) (Obj.), the CPU time in seconds (Time), and number of outer iterations (Iter.).
-We observe that all the compared algorithms attain comparable objective function value for different number of groups and regularization parameters $\mu$.
-In addition, the proposed MPGDA method consistently outperform the two RADA algorithms in terms of CPU time, which again validates its computational efficiency.
+(22, $0.1$) $2.187$ $0.408$ $265$ $2.187$ $0.451$ $154$ $2.187$ 0.106 $66$
+(44, $0.1$) $4.374$ $0.410$ $259$ $4.383$ $0.517$ $180$ $4.373$ 0.101 $64$
+(66, $0.1$) $6.560$ $0.400$ $224$ $6.569$ $0.643$ $223$ $6.560$ 0.103 $63$
+(88, $0.1$) $8.747$ $0.431$ $198$ $8.753$ $0.765$ $264$ $8.747$ 0.109 $66$
+($10$, $0.1$) $10.934$ $0.474$ $180$ $10.956$ $0.943$ $277$ $10.934$ 0.113 $64$
+(55, $0.1$) $5.467$ $0.416$ $244$ $5.494$ $0.719$ $240$ $5.467$ 0.110 $65$
+(55, $0.2$) $5.967$ $0.437$ $263$ $5.977$ $0.636$ $217$ $5.967$ 0.117 $71$
+(55, $0.5$) $7.469$ $0.467$ $286$ $7.489$ $0.672$ $229$ $7.467$ 0.155 $86$
+(55, $1.0$) $9.971$ $0.490$ $300$ $10.070$ $0.702$ $238$ $9.968$ 0.189 $97$
+Funding Qia Li is supported by the National Natural Science Foundation of China under grant 12471098, and by the Guangdong Province Key Laboratory of Computational National Science at Sun Yat-sen University Grant 2020B1212060032.
+Data Availability The default credit dataset used in Section 6.2 was downloaded from the website https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients.
+Declarations
+Conflict of interest  The authors declare that they have no conflict of interest.